### v1

# Trajectory Consistency Distillation

Abstract
Latent Consistency Model (LCM) extends the Consistency Model to the latent space and leverages the guided consistency distillation technique to achieve impressive performance in accelerating text-to-image synthesis. However, we observed that LCM struggles to generate images with both clarity and detailed intricacy. To address this limitation, we initially delve into and elucidate the underlying causes. Our investigation identifies that the primary issue stems from errors in three distinct areas. Consequently, we introduce Trajectory Consistency Distillation (TCD), which encompasses trajectory consistency function and strategic stochastic sampling. The trajectory consistency function diminishes the distillation errors by broadening the scope of the self-consistency boundary condition and endowing the TCD with the ability to accurately trace the entire trajectory of the Probability Flow ODE. Additionally, strategic stochastic sampling is specifically designed to circumvent the accumulated errors inherent in multi-step consistency sampling, which is meticulously tailored to complement the TCD model. Experiments demonstrate that TCD not only significantly enhances image quality at low NFEs but also yields more detailed results compared to the teacher model at high NFEs.
Diffusion, Consistency Models, Text-to-Image Generation
[…]
1 Introduction
Score-based generative models (SGMs), also commonly known as Diffusion Models (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Song et al., 2020b; Ho et al., 2020), have demonstrated their proficiency in various generative modeling domains such as image (Dhariwal & Nichol, 2021; Ramesh et al., 2022; Rombach et al., 2022), video (Ho et al., 2020; Wu et al., 2023; Guo et al., 2023), and audio (Kong et al., 2020; Chen et al., 2020; Popov et al., 2021), particularly in text-to-image synthesis (Nichol et al., 2022; Ramesh et al., 2022; Saharia et al., 2022; Podell et al., 2023). An noteworthy aspect of SGMs is the utilisation of stochastic differential equations (SDEs) and corresponding marginal-preserving ordinary differential equations (ODEs) to iteratively perturb data and eliminate noise (Song et al., 2020b). This facilitates an effective trade-off between generation cost and sampling quality, but they are also constrained by slow inference speed, requiring a substantial number of function evaluations (NFEs) to obtain satisfactory results.
To overcome this limitation, Song et al. (2023) proposed Consistency Models (CMs), an emerging class of powerful generative models capable of generating high-quality data with single-step or few-step sampling without the need for adversarial training. CMs are inextricably connected with SGMs in their underlying mathematical underpinnings, aiming to enforce the self-consistency property by mapping arbitrary points on the trajectory of the same Probability Flow Ordinary Differential Equation (PF ODE) to the trajectory’s origin (Song et al., 2020b). CMs can be trained with consistency distillation or treated as standalone generative models. Song et al. (2023) have demonstrated their superiority through extensive experiments in the pixel space. Latent Consistency Models (LCMs) (Luo et al., 2023a) further incorporate Latent Diffusion Models (LDMs) (Rombach et al., 2022) with CMs, achieving remarkable success in swiftly synthesizing high-resolution images conditioned on text. Moreover, LCM-LoRA (Luo et al., 2023b) improves the training efficiency of LCMs and converts them into a universal neural PF ODE solver by introducing LoRA (Hu et al., 2021) into the distillation process of LCMs. It is noteworthy that all these Consistency-Type Models still allow for striking a balance between computation and sample quality using Multistep Consistency Sampling (Song et al., 2023). In particular, allocating additional compute for more iterations can theoretically yield samples of higher quality.
Figure 1: Comparison between TCD and other state-of-the-art methods. TCD delivers exceptional results in terms of both quality and speed, completely surpassing LCM. Notably, LCM experiences a notable decline in quality at high NFEs. In contrast, TCD maintains superior generative quality at high NFEs, even exceeding the performance of DPM-Solver++(2S) with origin SDXL.
Despite the introduction of the guided distillation method and skipping-step technique (Luo et al., 2023a) by LCMs for effectively distilling knowledge from pre-trained diffusion models, the quality of images generated by LCMs in a single step or even with minimal steps (4$\sim$8) still lags significantly behind the convergence of its teacher model. Our investigation revealed that, in practice, increasing the number of inference iterations diminishes the visual complexity and quality of the results, as illustrated in Figure 1. This renders LCMs less capable of synthesizing samples that are perceptually comparable to those of their teacher model. Recent findings by EDM (Karras et al., 2022) identified that an excessive Langevin-like addition and removal of random noise result in a gradual loss of detail in the generated images. Additionally, Li et al. (2023) showed evidence that during the process of multi-step sampling, discrete errors accumulate over iterations, ultimately causing the generated images to deviate from the target distribution.
Drawing inspiration from the observation, we first meticulously examine the training procedure and delve into the multistep consistency sampling procedure to figure out the root cause. Our investigation revealed that the issue stems from the cumulative errors inherent in multi-step sampling processes. These errors predominantly originate from three sources: 1) the estimation errors in the original score matching model, 2) the distillation errors in the consistency model, and 3) the discretisation errors accumulated during the sampling phase. Collectively, these errors significantly undermine the efficacy of the multistep sampling consistency model, resulting in a performance that falls substantially short of expectations.
To suppress these errors, we introduce Trajectory Consistency Distillation (TCD), as summarized in Figure 2, which comprises two key elements: trajectory consistency function and strategic stochastic sampling. Specifically, inspired by the form of exponential integrators, the trajectory consistency function (TCF) diminishes distillation errors by expanding the boundary conditions of the consistency model and enabling seamless transitions at any point along the trajectory governed by the PF ODE, as demonstrated in Figure 1(a). Furthermore, strategic stochastic sampling (SSS) suppresses the accumulated discretisation error and estimation error according to the narrowed bi-directional iterations as shown in Figure 1(b).
Experiments show that TCD can significantly enhance the quality of images generated by LCM, surpassing it in performance. Furthermore, TCD is capable of outperforming the teacher model (e.g. SDXL with DPMSolver++) when sampling with sufficient iterations (e.g., 20 NFEs).
2 Preliminaries
2.1 Diffusion Models
Diffusion Models (DMs) start with a predefined forward process $\{\bm{x}_{t}\}_{t\in[0,T]}$ indexed by a continuous time variable $t$ with $T>0$, which progressively adds noise to data via Gaussian perturbations. The forward process can be modeled as a widely used stochastic differential equation (SDE) (Song et al., 2020b; Karras et al., 2022):
$\text{d}\bm{x}_{t}=\mu(t)\bm{x}_{t}\text{d}t+\nu(t)\text{d}\bm{w}_{t},$ (1)
where $\bm{w}_{t}$ denotes the $d$-dimensional standard Brownian motion and $\mu(t)\colon\mathbb{R}\rightarrow\mathbb{R}$ and $\nu(t)\colon\mathbb{R}\rightarrow\mathbb{R}$ are the drift and diffusion coefficients, respectively, where $d$ is the dimensionality of the dataset. Denote the marginal distribution of $\bm{x}_{t}$ following the forward process as $p_{t}(\bm{x}_{t})$ and, such an Itô SDE gradually perturbs the empirical data distribution $p_{0}(\bm{x})=p_{\rm{data}}(\bm{x})$ towards the prior distribution $p_{T}(\bm{x})\approx\pi(\bm{x})$ approximately, where $\pi(\bm{x})$ is a tractable Gaussian distribution.
(a) Training process, wherein the TCF expands the boundary conditions to an arbitrary timestep of $s$, thereby reducing the theoretical upper limit of error. The details can be found in Algorithm 1.
(b) Sampling process, as detailed in Algorithm 3 & Algorithm 4, where SSS significantly reduces accumulated error through the iterative traversal with the stochastic parameter $\gamma$ compared to the multistep consistency sampling.
Figure 2: The comparative overview of the baseline Consistency Distillation (Song et al., 2023) and the proposed Trajectory Consistency Distillation, includes Trajectory Consistency Function (TCF) for training and Strategic Stochastic Sampling (SSS) for inference.
Remarkably, Song et al. (2020b) proved that there exists an ordinary differential equation (ODE) dubbed the probability flow (PF) ODE, whose trajectories share the same marginal probability densities $\{p_{t}(\bm{x})\}_{t\in[0,T]}$ as the forward SDE,
$\frac{\text{d}\bm{x}_{t}}{\text{d}t}=\mu(t)\bm{x}_{t}-\frac{1}{2}\nu(t)^{2}\nabla_{\bm{x}}\log p_{t}(\bm{x}_{t}).$ (2)
As for sampling, the ground truth score in Equation 2 is approximated with the learned score model $\bm{s}_{\bm{\theta}}(\bm{x},t)\approx\nabla_{\bm{x}}\log p_{t}(\bm{x})$ via score matching (Hyvärinen et al., 2009; Song & Ermon, 2019; Ho et al., 2020). This yields an empirical estimate of the PF ODE, referred to as the empirical PF ODE:
$\frac{\text{d}\tilde{\bm{x}}_{t}}{\text{d}t}=\mu(t)\tilde{\bm{x}}_{t}-\frac{1}{2}\nu(t)^{2}\bm{s}_{\bm{\theta}}(\tilde{\bm{x}}_{t},t).$ (3)
Then samples can be drawn by solving the empirical PF ODE from $T$ to 0. There already exist off-the-shelf ODE solvers (Song et al., 2020b, a; Karras et al., 2022) or efficient numerical solvers (Lu et al., 2022a, b; Zhang & Chen, 2022) that can be directly applied to approximate the exact solution.
2.2 Consistency Models
Solving Equation 3 typically involves numerous neural network evaluations to generate decent samples. Thus, consistency models are proposed to directly map any points along the trajectory $\{\bm{x}_{t}\}_{t\in[0,T]}$ of the PF ODE to the origin of its trajectory, thereby facilitating generation in just a few steps. The associated mapping can be formuled as follows:
$\bm{f}(\bm{x}_{t},t)=\bm{x}_{0}\quad\forall t\in[0,T],$ (4)
with the boundary condition $\bm{f}(\bm{x}_{0},0)=\bm{x}_{0}$. It is worth noting that Equation 4 is equivalent to the self-consistency condition:
$\bm{f}(\bm{x}_{t},t)=\bm{f}(\bm{x}_{t}^{\prime},t^{\prime})\quad\forall t,t^{\prime}\in[0,T].$ (5)
A parametric model $\bm{f}_{\bm{\theta}}$ is constructed to estimate the consistency function $\bm{f}$ by enforcing the self-consistency property. Typically, $\bm{f}$ can be distilled from a pretrained diffusion model $F_{\bm{\theta}}(\bm{x}_{t},t)$ and parameterized as:
$\bm{f}_{\bm{\theta}}(\tilde{\bm{x}}_{t},t)=\begin{cases}\bm{x}_{0},&t=0\\ \texttt{Solver}(F_{\bm{\theta}}(\tilde{\bm{x}}_{t},t),t,0;\bm{\theta}),&t\in(0,T]\end{cases}$
The $\texttt{Solver}(\cdot,t,0;\theta)$ is the update function of an ODE solver used to estimate ${\bm{x}}_{0}$ when given the output $F_{\theta}(\tilde{\bm{x}}_{t},t)$ from a pretrained model at timestep $t$. For training consistency models, the objective of Consistency Distillation (CD) is defined as minimizing:
$\begin{split}\mathcal{L}_{\rm{CD}}^{N}&(\bm{\theta},\bm{\theta^{-}};\bm{\phi}):=\\ &\mathbb{E}\left[\lambda(t_{n})\left\lVert{\bm{f}_{\bm{\theta}}(\bm{x}_{t_{n+1}},t_{n+1})-\bm{f}_{\bm{{\theta^{-}}}}(\hat{\bm{x}}_{t_{n}}^{\bm{\phi}},t_{n})}\right\rVert_{2}^{2}\right],\end{split}$ (6)
where $0=t_{1}<t_{2}\cdots<t_{N}=T$, $n$ uniformly distributed over $\{1,2,\cdots,N-1\}$, $\lambda(\cdot)\in\mathbb{R}^{+}$ is a positive weighting function, and the expectation is taken with respect to $\bm{x}\sim p_{\rm{data}}$. $\bm{x}_{t_{n+1}}$ can be sampled using SDE  1 and $\hat{\bm{x}}_{t_{n}}^{\bm{\phi}}$ is calculated by $\hat{\bm{x}}_{t_{n}}^{\bm{\phi}}:=\Phi(\bm{x}_{t_{n+1}},t_{n+1},t_{n};\bm{\phi})$, where $\Phi(\cdots;\bm{\phi})$ represents the update function of a one-step ODE solver applied to the empirical PF ODE 3. Moreover, $\bm{\theta^{-}}$ is introduced to stabilize the training process and updated by an exponential moving average (EMA) strategy, i.e., given $10\leq\mu<$ 1, $\bm{\theta}^{-}\leftarrow\texttt{sg}(\mu\bm{\theta}^{-}+(1-\mu)\bm{\theta})$.
Besides the distillation strategy that needs an existing score model, Song et al. (2023) also introduced a way to train without any pre-trained models called Consistency Training (CT). In this paper, our primary focus is on the CD objective.
3 Elucidating Errors in Consistency Models
In this section, we elucidate various types of errors that occurred in multistep sampling of consistency models, which motivates us to propose corresponding solutions aimed at refining them. Our analysis reveals that the error is primarily composed of three components: distillation error of consistency models, estimation error of score matching models, and accumulated discretisation error (also known as truncation error) during solving the ODEs.
3.1 Consistency Distillation Error
With a well-trained model $\bm{\theta^{*}}$ such that $\mathcal{L}_{\rm{CD}}^{N}(\bm{\theta^{*}},\bm{\theta^{*}};\bm{\phi})=0$, and considering that the ODE solver $\Phi(\cdots;\bm{\bm{\phi}})$ has the local discretisation error, Theorem 1 in (Song et al., 2023) shows that the consistency distillation error is bounded:
$\mathop{\textnormal{sup}}\limits_{n,{\bm{x}}}\|{\bm{f}}_{\bm{\theta^{*}}}({\bm{x}},t_{n}),{\bm{f}}({\bm{x}},t_{n};\bm{\phi})\|_{2}=\mathcal{O}\left((\Delta t)^{p}\right),$ (7)
with $\Delta t$ and $p$ as defined in (Song et al., 2023).
3.2 Error Bound in Multistep Consistency Sampling
Theoretically, with a well-trained consistency model ${\bm{f}}_{\bm{\theta^{*}}}$, one can generate samples with just one forward pass through the consistency model. However, the one-step sampling yields suboptimal results (Luo et al., 2023a; Song et al., 2023). Thus, multistep consistency sampling has been introduced in (Song et al., 2023) to enhance sample quality through alternating denoising and noise injection steps. We term it as multistep sampling in this manuscript.
For simplicity, we consider VP SDEs in this paper; thus, the drift and diffusion coefficients $\mu(t)$ and $\nu(t)$ in SDE( Equation 1) can be written as:
$\mu(t)=\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t},\quad\nu(t)=\sqrt{\frac{\mathrm{d}\sigma_{t}^{2}}{\mathrm{d}t}-2\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}\sigma_{t}^{2}},$ (8)
where $\alpha_{t}$ and $\sigma_{t}$ specify the noise schedule in the perturbation kernels,
$q_{0t}({\bm{x}}_{t}|{\bm{x}}_{0})=\mathcal{N}({\bm{x}}_{t}|\alpha_{t}{\bm{x}}_{0},\sigma_{t}^{2}\bm{I}).$ (9)
Given a sequence of $N$ sampling timesteps $T=\tau_{1}>\tau_{2}>\cdots>\tau_{N}$ and an initial value $\tilde{{\bm{x}}}_{\tau_{1}}\sim\mathcal{N}(0,\bm{I})$, the generating procedure of the $n$-th step can be written as:
${\bm{x}}_{\tau_{1}\rightarrow 0}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}(\tilde{{\bm{x}}}_{\tau_{1}},T),$ (10)
Diffuse: $\tilde{{\bm{x}}}_{\tau_{n}}\leftarrow\alpha_{\tau_{n}}{\bm{x}}_{\tau_{(n-1)}\rightarrow 0}+\sigma_{\tau_{n}}{\bm{z}},\quad{\bm{z}}\sim\mathcal{N}(0,\bm{I})$
Denoise: ${\bm{x}}_{\tau_{n}\rightarrow 0}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}(\tilde{{\bm{x}}}_{\tau_{n}},\tau_{n}),$
this process is detailed in Algorithm 3. We derive a further corollary from Corollary 7 in  (Lyu et al., 2023) as follows:
Corollary 3.1.
With the sampling process defined in Equation 10 and denoting the distribution of ${\bm{x}}_{\tau_{n}\rightarrow 0}$ as $q_{\bm{\theta^{*}},n}$, we have single-step sampling result, $q_{\bm{\theta^{*}},1}={\bm{f}}_{{\bm{\theta^{*}}},T}\sharp\mathcal{N}(0,\bm{I})$ and multistep sampling result, $q_{\bm{\theta^{*}},N}={\bm{f}}_{{\bm{\theta^{*}}},\tau_{N}}\sharp\mathcal{N}(\alpha_{\tau_{N}}{\bm{x}}_{\tau_{(N-1)}\rightarrow 0},\sigma_{\tau_{N}}\bm{I})$ with $N>1$. The Total Variational (TV) distance between $q$ and $p_{\rm{data}}$ is
$TV(q_{\bm{\theta^{*}},1},p_{\rm{data}})$ $=\mathcal{O}\left(T(\varepsilon_{cd}+\mathcal{L}_{f}\varepsilon_{se})\right)$
$TV(q_{\bm{\theta^{*}},N},p_{\rm{data}})$ $=\mathcal{O}\left({2^{-N}T(\varepsilon_{cd}+\mathcal{L}_{f}\varepsilon_{se})}\right),$
where normal-$\sharp$ is a push-forward operator associated with a measurable map, $\mathcal{L}_{f}$ is the Lipschitz constant of consistency model, and $\varepsilon_{cd}$, $\varepsilon_{se}$ represent the error of consistency distillation in Equation 7 and score matching estimation, respectively.
The detailed proof can be found in Appendix D.1. From  Corollary 3.1, we observe that multistep sampling has a lower error bound than a single step. This observation aligns with the empirical findings presented in (Karras et al., 2022): the local truncation error scales superlinearly with respect to step size, thus increasing $N$ improves the accuracy of the solution. Moreover, it illustrates that the model performance tends to underperform single-step generation methods (Sauer et al., 2023; Yin et al., 2023) without the help of additional supervision.
3.3 Accumulated Error in Multistep Sampling
However, due to the error of estimation and discretisation in practice, the total error will be accumulated in every sampling step. Denote the distribution of ${\bm{x}}_{\tau_{n}\rightarrow 0}$ output by the true consistency function ${\bm{f}}(\cdot,\cdot;\bm{\phi})$ as $p_{n}$, Theorem 2 in (Chen et al., 2022) demonstrates that the TV error between $q_{\bm{\theta^{*}},n}$ and $p_{n}$ is
$TV(q_{\bm{\theta^{*}},n},p_{n})=\mathcal{O}(\sqrt{\tau_{n}}),$ (11)
the error in each $n$-th step will accumulate over $N$ sampling steps, resulting in the defined accumulated error
$TV(q_{\bm{\theta^{*}},N},p_{N})=\mathcal{O}\left(\sum_{n=1}^{N}\sqrt{\tau_{n}}\right).$ (12)
The above result is a special case of Theorem 4.2, and its proof is presented in Section D.3. As a consequence, there is a significant reduction in image details at higher NFEs, as shown in Figure 3.
4 Trajectory Consistency Distillation
4.1 Trajectory Consistency Function
Definition.
The distillation error from the consistency model can be reduced by expanding the original boundary conditions to encompass the entire trajectory. To facilitate this, we introduce the Trajectory Consistency Function (TCF), designed to enable comprehensive tracking along the full trajectory.
${\bm{f}}({\bm{x}}_{t},t,s)\mapsto{\bm{x}}_{s}.$ (13)
The trajectory consistency function possesses the property of trajectory consistency, enhancing the original self-consistency property in an endpoint-unrestricted trajectory. Specifically, its output remains consistent for arbitrary sets $({\bm{x}}_{t},t)$ with the given $s$ that belong to the same PF ODE trajectory, where $0\leqslant s\leqslant t\leqslant T$,
${\bm{f}}({\bm{x}}_{t},t,s)={\bm{f}}({\bm{x}}^{\prime}_{t},t^{\prime},s)\quad\forall t,t^{\prime},s\in[0,T].$ (14)
Figure 3: Synthesis results across various NFEs. Due to accumulated errors in multistep sampling, LCM experiences a loss of image detail, leading to a degradation in performance, whereas TCD addresses this issue. Additional samples are available in Section E.1.
Parameterisation.
The semi-linear structure of the empirical PF-ODE revealed by (Lu et al., 2022a, b; Zhang & Chen, 2022) motivates us to parameterise the trajectory consistency function using the exponential integrators form, as shown in Equation 15.
${\bm{f}}_{\bm{\theta}}({\bm{x}}_{t},t,s)=\frac{\sigma_{s}}{\sigma_{t}}{\bm{x}}_{t}+\sigma_{s}\int_{\lambda_{t}}^{\lambda_{s}}e^{\lambda}\hat{{\bm{x}}}_{\bm{\theta}}(\hat{{\bm{x}}}_{\lambda},\lambda)\text{d}\lambda,$ (15)
where $\lambda_{t}:=\log(\alpha_{t}/\sigma_{t})$ is the log-SNR, and ${\bm{x}}_{\bm{\theta}}$ is a trainable network with parameter $\bm{\theta}$.
For $k\geqslant 1$, one can take the ($k$-1)-th Taylor expansion at $\lambda_{t}$ for ${\bm{x}}_{\bm{\theta}}$ w.r.t $\lambda\in[\lambda_{s},\lambda_{t}]$, we have:
$\begin{split}{\bm{f}}_{\bm{\theta}}&({\bm{x}}_{t},t,s)=\frac{\sigma_{s}}{\sigma_{t}}{\bm{x}}_{t}+\\ &\sigma_{s}\sum_{n=0}^{k-1}{\bm{x}}_{\bm{\theta}}^{(n)}(\hat{{\bm{x}}}_{\lambda_{t}},\lambda_{t})\int_{\lambda_{t}}^{\lambda_{s}}e^{\lambda}\frac{(\lambda-\lambda_{t})^{n}}{n!}\text{d}\lambda+\mathcal{O}(h^{k+1}),\end{split}$ (16)
where $h=\lambda_{s}-\lambda_{t}$ and ${\bm{x}}_{\bm{\theta}}^{(n)}(\cdot,\cdot)$ is the $n$-th order total derivatives of ${\bm{x}}_{\bm{\theta}}$ w.r.t $\lambda$.
Here, we consider 1st-order and 2nd-order estimations, omitting high-order terms $\mathcal{O}(h^{k+1})$.
[…]
For the $2$nd-order expansion, we can write the trajectory consistency function as:
$\begin{split}\text{TCF(2):}\quad{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t},&t,s)=\frac{\sigma_{s}}{\sigma_{t}}{\bm{x}}_{t}-\alpha_{s}(e^{-h}-1)\\ &\left((1-\frac{1}{2r})\hat{\bm{x}}_{\theta}({\bm{x}}_{t},t)+\frac{1}{2r}\hat{\bm{x}}_{\theta}(\hat{\bm{x}}_{u},u)\right),\end{split}$ (18)
where $u$ is the mid-timestep w.r.t. $t>u>s$ and $r:=(\lambda_{u}-\lambda_{t})/h$.
We also propose a modified network $F_{\bm{\theta}}$ with additional parameters for conditioning on $s$ to directly estimate the exponentially weighted integral of ${\bm{x}}_{\bm{\theta}}$ without omitting residual term when $k=1$ in Equation 16:
[…]
Broadening the Boundary Condition.
The boundary condition in the CM is limited to intervals extending from any start point on the solution trajectory to the origin. Conversely, TCF mitigates this constraint, allowing the model to handle any interval along the PF-ODE trajectory and preventing trivial solutions ${\bm{f}}_{\bm{\theta}}({\bm{x}}_{t},t,s)\equiv 0$ from arising in TCD training. Consequently, we broaden the boundary condition to encompass a more comprehensive range of trajectory intervals,
${\bm{f}}_{\bm{\theta}}({\bm{x}}_{s},s,s)={\bm{x}}_{s},$ (20)
it is evident that all three types of our parametrisation satisfy the broad boundary condition effortlessly.
Training.
Considering $0=t_{1}<t_{2}\cdots<t_{N}=T$ and given the one-step update function of a trained PF ODE solution $\Phi(\cdots;\bm{\phi})$ parameterised by $\bm{\phi}$, we can obtain an accurate estimation ${\bm{x}}_{t_{n}}$ from ${\bm{x}}_{t_{n+k}}$ by executing $k$ discretisation steps with $\Phi^{(k)}(\cdots;\bm{\phi})$,
$\hat{{\bm{x}}}_{t_{n}}^{\phi,k}=\Phi^{(k)}({\bm{x}}_{t_{n+k}},t_{n+k},t_{n};\bm{\phi}).$ (21)
Thus, we could express the object of trajectory distillation in alignment with reconstruction:
$\mathcal{L}^{N}_{\rm{TCD}}(\bm{\theta},$ $\bm{\theta^{-}};\bm{\phi}):=\mathbb{E}[\omega(t_{n},t_{m})$ (22)
$\|{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n+k}},t_{n+k},t_{m})-{\bm{f}}_{\bm{\theta}^{-}}(\hat{{\bm{x}}}_{t_{n}}^{\phi,k},t_{n},t_{m})\|^{2}_{2}],$
where $n\sim\mathcal{U}[\![1,N-1]\!]$, $m\sim\mathcal{U}[\![1,n]\!]$, $\bm{\theta}^{-}$ can be either updated by EMA: $\bm{\theta}^{-}\leftarrow\texttt{sg}(\mu\bm{\theta}^{-}+(1-\mu)\bm{\theta})$ or stop the gradient without updating: $\texttt{sg}(\bm{\theta})$, $\omega(\cdots)$ is a positive weighting function, and we find $\omega(t_{n},t_{m})\equiv 1$ performs well in our experiments. We also employ the skipping-step method proposed in (Luo et al., 2023a) to accelerate convergence. The detailed training process is outlined in Algorithm 1.
Below, we offer a theoretical justification based on asymptotic analysis to explain how trajectory consistency distillation optimizes the distillation error.
Theorem 4.1.
Let $\Delta t:=\rm{max}_{n\in[\![1,N-1]\!]}\{\lvert t_{n+1}-t_{n}\rvert\}$, and $\bm{f}(\cdot,\cdot,\cdot;\bm{\phi})$ be the trajectory consistency function of the empirical PF ODE in Equation 3. Assume $\bm{f}_{\bm{\theta}}$ satisfies the Lipschitz condition, that is, there exists $L>0$ such that for all $t\in[0,T]$, $\bm{x}$ and $\bm{y}$, we have $\left\lVert{\bm{f}_{\bm{\theta}}(\bm{x},t,s)-\bm{f}_{\bm{\theta}}(\bm{y},t,s)}\right\rVert_{2}\leqslant L\left\lVert{\bm{x}-\bm{y}}\right\rVert_{2}$. Assume further that for all $n\in[\![1,N-1]\!]$, the $p$-th order ODE solver called at $t_{n+1}$ has local error uniformly bounded by $\mathcal{O}((t_{n+1}-t_{n})^{p+1})$ with $p\geqslant 1$. Then, if there is a $\bm{\theta^{*}}$ so that $\mathcal{L}_{\rm{TCD}}^{N}(\bm{\theta^{*}},\bm{\theta^{*}};\bm{\phi})=0$, for any $n\in[\![1,N-1]\!]$ and $m\in[\![1,n]\!]$, we have
$\begin{split}\mathop{\textnormal{sup}}\limits_{n,m,{\bm{x}}}\|{\bm{f}}_{\bm{\theta^{*}}}({\bm{x}},t_{n},t_{m}),{\bm{f}}({\bm{x}},&t_{n},t_{m};\phi)\|_{2}\\ &=\mathcal{O}\left((\Delta t)^{p}\right)(t_{n}-t_{m}).\end{split}$
Proof.
The proof is provided in  Section D.2. ∎
Theorem 4.1 implies that the distillation error of TCF is upper bound by that of CD presented in Equation 7.
4.2 Strategic Stochastic Sampling
The proposed trajectory consistency function not only optimizes the distillation loss but also enables the model to access non-origin destinations along the PF ODE. This capability allows Strategic Stochastic Sampling (SSS) to further reduce discretisation errors and estimation errors introduced in each sampling step.
Specifically, every sampling step in SSS includes the denoise sub-step according to the ODE solver and the diffuse sub-step based on Langevin SDE. In comparison with multistep consistency sampling, where the endpoint and noise level are fixed, SSS introduces the additional parameter $\gamma$ to control the destination point for the denoise step and allows for the adjustment of random noise level for the diffuse step, as detailed in Figure 1(b) and Algorithm 4. This parameter is referred to as the stochastic parameter.
In the denoising sub-step, our focus is on reducing the errors that arise from the $n$-th step in Equation 11 to $\mathcal{O}(\sqrt{\tau_{n}-(1-\gamma)\tau_{(n+1)}})$ when $n\in[\![1,N-1]\!]$, by the prediction of the non-origin, and ultimately optimize the accumulated error as shown in Theorem 4.2.
Theorem 4.2.
As the strategic stochastic sampling process defined in Algorithm 4, and denoting the $N>1$ steps sampling results with trained trajectory consistency model ${\bm{f}}_{\bm{\theta^{*}}}$ as $q_{\bm{\theta^{*}},N}={\bm{f}}_{{\bm{\theta^{*}}},\tau_{N}}\sharp\mathcal{N}(\alpha_{\tau_{N}}{\bm{x}}_{\tau_{(N-1)}\rightarrow 0},\sigma_{\tau_{N}}\bm{I})$, the results with exact trajectory consistency function ${\bm{f}}_{\bm{\theta}}$ as $p_{N}$, then the accumulated error in SSS is defined as:
$TV(q_{\bm{\theta^{*}},N},p_{N})=\mathcal{O}\left(\sum_{n=1}^{N-1}\sqrt{\tau_{n}-(1-\gamma)\tau_{n+1}}+\sqrt{\tau_{N}}\right),$ (23)
where $\gamma\in[0,1]$ is the parameter controlling the destination in each denoising step.
Proof.
The proof is provided in Section D.3. ∎
While sufficient stochasticity in diffuse sub-step helps reduce both the estimation errors and discretisation accumulated in earlier sampling steps and drives the sample towards the desired marginal distribution, a similar finding is also observed in (Karras et al., 2022; Xu et al., 2023b). In SSS, we encourage the step size of the diffuse sub-step to be smaller than that of the denoise sub-step, which is contrary to (Xu et al., 2023b). It is also worth noting that when $\gamma$ is low, the estimation error plays a more important role. Hence, the optimal value of $\gamma$ should be determined empirically, as we show in Section 5.3.
Figure 4: Qualitative comparison. For each prompt, images are generated using the same random seed for every model, without any cherry-picking. More results are provided in Section E.3.
Figure 5: Qualitative effects of stochastic parameter $\gamma$ with same NFEs. Images under the same prompt but with different $\gamma$ applied during sampling. The leftmost image is sampled from LCM (Luo et al., 2023b). More samples can be found in Section E.2.
Table 1: Quantitative comparison on the COCO validation set.
[…]
  • The best scores are highlighted in bold, and the runner-ups are underlined.
Figure 6: Quantitative ablation on different stochastic parameter $\gamma$.
4.3 Extension to Large Text Conditional Models
Conditional models frequently outperform their unconditional counterparts and exhibit a wider range of use cases (Bao et al., 2022; Dhariwal & Nichol, 2021; Ho & Salimans, 2022). Particularly, text conditional models have recently garnered substantial attention, showcasing remarkable results (Nichol et al., 2022; Ramesh et al., 2022; Saharia et al., 2022; Podell et al., 2023). The trajectory consistency function can be seamlessly integrated into conditional models by introducing an additional input, ${\bm{c}}$, to accommodate conditioning information, such as text. This results in the transformation of the trajectory function to ${\bm{f}}_{\rm{\theta}}({\bm{x}}_{t},{\bm{c}},t,s)$, and the guided distillation method proposed by (Meng et al., 2023; Luo et al., 2023a) can be directly integrated into it, as detailed in Algorithm 2.
Trajectory consistency distillation can directly occur on top of the parameters of a pre-trained diffusion model as the fine-tuning process. For scaling TCD to larger models (e.g., SDXL) with significantly reduced memory consumption, we incorporate Low-Rank Adaptation (LoRA) (Hu et al., 2021), a parameter-efficient fine-tuning method, into the distillation process. Additionally, the parameters of LoRA can be identified as a versatile acceleration module applicable to different fine-tuned models or LoRAs sharing the same base model without the need for additional training, aligning with the observations in (Luo et al., 2023b).
5 Experiments
5.1 Experimental Setup
We selected SDXL (Podell et al., 2023), a widely recognized diffusion model, as our backbone. By default, we employ TCF(1) as the parameterisation and set the stochastic parameter $\gamma$ as 0.2. The influence of $\gamma$ and parameterisation type is left to be explored in the ablation studies (Section 5.3). For detailed implementation information, please refer to Appendix C.
Figure 7: Versatility of TCD. Our TCD LoRA can be directly applied to a wide range of models, including custom community models, styled LoRA, ControlNet, and IP-Adapter, accelerating their generation in just a few steps with high quality.
Table 2: Quantitative ablation on the TCF parameterisation type.
Para Type FID $\downarrow$ IC Score $\uparrow$ ImageReward $\uparrow$ PickScore $\uparrow$
TCF(1) 12.68 0.5095 68.49 22.31
TCF(2) 13.35 0.5037 58.13 22.07
TCF(S+) 13.03 0.4176 57.96 22.01
5.2 Main Results
To illustrate the effectiveness and superiority of our methods, we perform qualitative and quantitative comparisons with prior works, including Euler (Karras et al., 2022), efficient numerical ODE solvers like DDIM (Song et al., 2020a) and DPM-Solver++(2S) (Lu et al., 2022b), and a relevant work LCM (Luo et al., 2023a) as baseline.
Qualitative Results.
As illustrated in  Figure 4, prior efficient numerical methods produce suboptimal images with 4 steps, while LCM can generate relatively better images. Our TCD further enhances visual quality. With an increased number of function evaluations (20 steps), the quality of samples generated by DDIM or DPM-Solver++(2S) improves rapidly. However, the improvement in LCM is not as noticeable, resulting in smoother and less detailed images due to the accumulated error in multistep sampling. In contrast, TCD addresses this flaw, producing more detailed images that surpass even the teacher model, e.g., SDXL with DDIM.
Quantitative Comparison.
For the quantitative evaluation, we conducted zero-shot image generation using 5K captions from COCO2017 validation set. We employed the Frechet Inception Distance (FID) and the Image Complexity Score (Feng et al., 2022) to assess the sample quality and visual complexity of generated images. Additionally, we used Image Reward (Xu et al., 2023a) and PickScore (Kirstain et al., 2023) to quantify text-image alignment and human preference. As depicted in Table 1, TCD shows an improvement in the performance of LCM across various timesteps and metrics. In comparison to numerical ODE-Solver, TCD exhibits the ability to generate high-quality images in only 2 or 4 steps, surpassing them significantly. When increasing steps to 8 or 20, LCM’s performance tends to decline, while TCD even outperforms its teacher in almost all metrics.
5.3 Ablation Studies
Effects of Stochastic Parameter.
The effect of the stochastic parameter $\gamma$ is shown in Figure 5 and Figure 6. As illustrated in Figure 5 and  Figure 6 (b), we find that with an increase in $\gamma$, the visual complexity and fineness of the image gradually improve. Furthermore, our observation indicates that implementing a stochastic sampling process can reduce accumultaed errors. Notably, when the gamma value is set to 0, the inherent estimation error of the model becomes more pronounced, leading to a decrease in quality evaluated with FID, as shown in  Figure 6 (a).
Parameterisation Type.
We illustrate the impact of our proposed different parameterisation types in Table 2 with 4 sampling steps. We observed that the instability of the high-order solver discussed in (Lu et al., 2022b) also exists in TCF(2), resulting in slightly inferior performance compared to TCF(1). Due to the introduction of additional parameters, TCF(S+) faces challenges in effectively distilling the teacher’s information, leading to poorer performance.
Versatility of TCD.
To assess the versatility of TCD, we extensively tested it on various models, including the popular community model Animagine XL V3  [^1] Animagine: https://civitai.com/models/260267/animagine-xl-v3, styled LoRA Papercut XL  [^2] Papercut: https://civitai.com/models/122567/papercut-sdxl, Depth ControlNet  [^3] Depth ControlNet: https://huggingface.co/diffusers/controlnet-depth-sdxl-1.0, Canny ControlNet  [^4] Canny ControlNet: https://huggingface.co/diffusers/controlnet-canny-sdxl-1.0, and IP-Adapter  [^5] IP-Adapter: https://github.com/tencent-ailab/IP-Adapter. The results shown in Figure 7 imply that TCD can be directly applied to various models to accelerate image generation with high quality in only 2-8 steps. Additional samples can be found in Section E.4.
6 Conclusion
In this work, we introduce TCD, a novel distillation method that reduces inherent errors present in consistency models, including TCF for training and SSS for sampling. The TCF is proposed to diminish distillation errors and enable the model to track the trajectory along the PF ODE. Furthermore, SSS is proposed to reduce accumulated errors by the bijective traversal. Remarkably, TCD outperforms LCM across all sampling steps and exhibits superior performance compared to numerical methods of teacher model. We believe that TCD can provide novel perspectives for fast and high-quality image generation, while certain characters of TCD also contribute valuable insights to downstream applications, e.g., enhanced details for super-resolution and a better intermediate manifold for editing.
Limitations.
In our experiments, we observed instability in high-order TCF and poor convergence in TCF(S+). Further analysis is necessary to ascertain the stability of the high-order function and TCF(S+). Additionally, it is worth investigating an improved design to achieve fewer steps generation, e.g., single step.
Impact Statements
While our advancements in sample quality and speed can help reduce inference costs, they also have the potential to amplify negative societal effects, such as disseminating disinformation. In the future, it is encouraged to implement adequate guardrails and detection techniques to minimize the risk of misuse.
Appendix A Related Works
Diffusion Models.
Diffusion Models (DMs), also known as Score-based generative models, have shown superior performance in various generative fields. Chen et al. (2022) provides theoretical convergence guarantees, implying that DMs can efficiently sample from essentially any realistic data distribution under minimal data assumptions. ADM (Dhariwal & Nichol, 2021) firstly shows the potential of DMs to outperform GANs. EDM (Karras et al., 2022) further elucidates the design space of GMs, clearly separating the concrete design choices, deriving best practices for the sampling process, and improving the training dynamics, thereby drastically improving the results.
Text-conditional Diffusion Models.
DMs have specially achieved great success in the area of text-to-image synthesis (Nichol et al., 2022; Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022; Balaji et al., 2022; Podell et al., 2023). To reduce computational cost, diffusion models typically operate within the latent space (Rombach et al., 2022; Podell et al., 2023) or include separate super-resolution steps (Ramesh et al., 2022; Saharia et al., 2022; Balaji et al., 2022). The integration of classifier-free diffusion guidance (Ho & Salimans, 2022; Dieleman, 2022) during the sampling process dramatically improves samples produced by conditional diffusion models at almost no additional cost.
Fast Sampling of DMs.
DMs exhibit great generative capabilities but are bottlenecked by their slow sampling speed. Various numerical methods driven by the SDE mathematical underpinnings underlying DMs have been proposed for accelerating DM sampling. DDIM (Song et al., 2020a) has originally shown promise for few-step sampling. Other works involve predictor-corrector samplers (Song et al., 2020b; Karras et al., 2022), exponential integrators (Lu et al., 2022a; Zhang & Chen, 2022; Lu et al., 2022b), and automated methods to adjust solvers (Kong & Ping, 2021; Jolicoeur-Martineau et al., 2021). Despite these methods having achieved great improvement in fast sampling, they are limited by the inherent discretisation error present in all solvers (De Bortoli et al., 2021). Ultimately, the sample quality obtained with few NFEs is limited. Another series of research works exemplified by distillation techniques (Luhman & Luhman, 2021; Salimans & Ho, 2022; Meng et al., 2023). These methods distill knowledge from pretrained models into few-step samplers, representing an efficient solution for few NFEs. However, they may experience a lengthy and costly process that requires huge amounts of data and suffer from slow convergence.
Consistency Models.
To overcome the limitations of current fast samplers, Song et al. (2023) proposes Consistency Models (CMs) learning a direct mapping from noise to data built on the top of the trajectory of PF ODE, achieving one-step generation while allowing multi-step sampling to trade compute for quality. Lyu et al. (2023) provide the first convergence guarantees for CMs under moderate assumptions. Kim et al. (2023) proposes a universal framework for CMs and DMs. The core design is similar to ours, with the main differences being that we focus on reducing error in CMs, subtly leverage the semi-linear structure of the PF ODE for parameterization, and avoid the need for adversarial training. Latent Consistency Models (LCMs) (Luo et al., 2023a) integrate consistency distillation with latent diffusion models (Rombach et al., 2022) to achieve impressive performance in accelerating text-to-image synthesis. LCM-LoRA (Luo et al., 2023b) further improves the training efficiency and versatility of LCMs by introducing LoRA (Hu et al., 2021) into the distillation process. However, due to inherent accumulated errors in multistep consistency samplings, LCM experiences a notable decline in quality at high NFEs. In contrast, our TCD leverages trajectory consistency and bijective traversal to eliminate this defect.
Parameter-Efficient Fine-Tuning.
Training a diffusion model is highly resource-intensive and environmentally unfriendly. Fine-tuning such a model can also be challenging due to the vast number of parameters involved (Aghajanyan et al., 2020). Thus, Parameter-Efficient Fine-Tuning (PEFT) (Houlsby et al., 2019) was proposed to enable the fine-tuning of pretrained models with a limited number of parameters required for training. Among these techniques, Low-Rank Adaptation (LoRA) (Hu et al., 2021) has demonstrated transcendental performance. LoRA’s strategy involves freezing the pretrained model weights and injecting trainable rank decomposition matrices, which succinctly represent the required adjustments in the model’s weights for fine-tuning. With this strategy, LoRA significantly reduces the volume of parameters to be modified, thereby substantially decreasing both computational load and storage demands.
Appendix B Algorithmic Details
B.1 Algorithm Details of Trajectory Consistency Distillation
Due to space limitations, we omitted some implementation details in the main body, but we provided a detailed algorithm for trajectory consistency distillation during training as Algorithm 1 and a guided version for conditional models as Algorithm 2.
Algorithm 1 Trajectory Consistency Distillation
  Input: dataset $\mathcal{D}$, initial model parameter $\bm{\theta}$, learning rate $\eta$, the update function of ODE solver $\Phi(\cdots;\bm{\phi})$, EMA rate $\mu$, noise schedule $\alpha_{t}$, $\sigma_{t}$, skipping interval $k$ $\bm{\theta}^{-}\leftarrow\bm{\theta}$
  repeat
     Sample ${\bm{x}}\sim\mathcal{D}$
     Sample $n\sim\mathcal{U}[\![1,N-k]\!]$, $m\sim\mathcal{U}[\![1,n]\!]$
     Sample ${\bm{x}}_{t_{n+k}}\sim\mathcal{N}(\alpha_{t_{n+k}}{\bm{x}},\sigma_{t_{n+k}}^{2}\bm{I})$
     $\hat{{\bm{x}}}_{t_{n}}^{\phi,k}\leftarrow\Phi^{(k)}({\bm{x}}_{t_{n+k}},t_{n+k},t_{n};\bm{\phi})$
     $\mathcal{L}(\bm{\theta},\bm{\theta^{-}};\bm{\phi})\leftarrow$ $\|{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n+k}},t_{n+k},t_{m})-{\bm{f}}_{\bm{\theta}^{-}}(\hat{{\bm{x}}}_{t_{n}}^{\phi,k},t_{n},t_{m})\|^{2}_{2}$
     $\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta},\bm{\theta^{-}};\bm{\phi})$
     $\bm{\theta^{-}}\leftarrow\texttt{sg}(\mu\bm{\theta}^{-}+(1-\mu)\bm{\theta})$
  until convergence
Algorithm 2 Guided Trajectory Consistency Distillation
  Input: dataset $\mathcal{D}$, initial model parameter $\bm{\theta}$, learning rate $\eta$, the update function of ODE solver $\Phi(\cdots;\bm{\phi})$, EMA rate $\mu$, noise schedule $\alpha_{t}$, $\sigma_{t}$, skipping interval $k$ $\bm{\theta}^{-}\leftarrow\bm{\theta}$
  repeat
     Sample $({\bm{x}},{\bm{c}})\sim\mathcal{D}$, $\omega\sim\mathcal{U}[\omega_{\text{min}},\omega_{\text{max}}]$
     Sample $n\sim\mathcal{U}[\![1,N-k]\!]$, $m\sim\mathcal{U}[\![1,n]\!]$
     Sample ${\bm{x}}_{t_{n+k}}\sim\mathcal{N}(\alpha_{t_{n+k}}{\bm{x}},\sigma_{t_{n+k}}^{2}\bm{I})$
     $\hat{{\bm{x}}}_{t_{n}}^{\phi,k}\leftarrow(1+\omega)\Phi^{(k)}({\bm{x}}_{t_{n+k}},t_{n+k},t_{n},{\bm{c}};\bm{\phi})-$ $\omega\Phi^{(k)}({\bm{x}}_{t_{n+k}},t_{n+k},t_{n},{\bm{c}};\bm{\phi})$
     $\mathcal{L}(\bm{\theta},\bm{\theta^{-}};\bm{\phi})\leftarrow$ $\|{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n+k}},{\bm{c}},t_{n+k},t_{m})-{\bm{f}}_{\bm{\theta}^{-}}(\hat{{\bm{x}}}_{t_{n}}^{\phi,k},{\bm{c}},t_{n},t_{m})\|^{2}_{2}$
     $\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta},\bm{\theta^{-}};\bm{\phi})$
     $\bm{\theta^{-}}\leftarrow\texttt{sg}(\mu\bm{\theta}^{-}+(1-\mu)\bm{\theta})$
  until convergence
B.2 The Difference between Multistep Consistency Sampling and Strategic Stochastic Sampling
For simplicity, we consider VP SDEs in this paper. Thus, the perturbation kernels satisfy:
$q_{0t}({\bm{x}}_{t}|{\bm{x}}_{0})=\mathcal{N}({\bm{x}}_{t}|\alpha_{t}{\bm{x}}_{0},\sigma_{t}^{2}\bm{I}),$ (24)
where $\alpha_{t}$ and $\sigma_{t}$ specify the noise schedule. While the transition kernels for VP SDE have the form:
$q({\bm{x}}_{t}|{\bm{x}}_{t-1})=\mathcal{N}\left(\frac{\alpha_{t}}{\alpha_{t-1}}{\bm{x}}_{t-1},\sqrt{1-\frac{\alpha_{t}^{2}}{\alpha_{t-1}^{2}}}\bm{I}\right).$ (25)
The detailed implementation of multistep consistency sampling in (Song et al., 2023) and our proposed strategic stochastic sampling are presented in Algorithm 3 and Algorithm 4, respectively. Both of them alternate between a denoising sub-step and a diffusing sub-step. The key distinction lies in the fact that the ending point of multistep consistency sampling is always set to 0. In contrast, strategic stochastic sampling introduces control through the parameter $\gamma$, allowing for the adjustment of the random noise level during alternating denoise and diffuse steps.
Algorithm 3 Multistep Consistency Sampling
  Input: Consistency model ${\bm{f}}_{\bm{\theta^{*}}}(\cdot,t)$, sequence of time points $T=\tau_{1}>\tau_{2}>\cdots>\tau_{N}$, initial value $\tilde{{\bm{x}}}_{\tau_{1}}=\tilde{{\bm{x}}}_{T}\sim\mathcal{N}(0,\bm{I})$
  for $n=1$ to $N-1$ do
     Sample ${\bm{z}}\sim\mathcal{N}(0,\bm{I})$
     Fixed ending time $\tau_{n}^{{}^{\prime}}=0$ Denoise
     ${\bm{x}}_{\tau_{n}\rightarrow\tau_{n}^{{}^{\prime}}}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}(\tilde{{\bm{x}}}_{\tau_{n}},\tau_{n})$ Diffuse
     $\tilde{{\bm{x}}}_{\tau_{(n+1)}}\leftarrow\alpha_{\tau_{(n+1)}}{\bm{x}}_{\tau_{n}\rightarrow\tau_{n}^{{}^{\prime}}}+\sigma_{\tau_{(n+1)}}{\bm{z}}$
  end for
  ${\bm{x}}_{\tau_{N}\rightarrow 0}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}(\tilde{{\bm{x}}}_{\tau_{N}},\tau_{N})$
  Output: ${\bm{x}}_{\tau_{N}\rightarrow 0}$
Algorithm 4 Strategic Stochastic Sampling
  Input: Trajectory consistency model ${\bm{f}}_{\bm{\theta^{*}}}(\cdot,t,s)$, sequence of time points $T=\tau_{1}>\tau_{2}>\cdots>\tau_{N}$, initial value $\tilde{{\bm{x}}}_{\tau_{1}}=\tilde{{\bm{x}}}_{T}\sim\mathcal{N}(0,\bm{I})$, parameter $\gamma\in[0,1]$
  for $n=1$ to $N-1$ do
     Sample ${\bm{z}}\sim\mathcal{N}(0,\bm{I})$
     Calculate ending time $\tau_{n}^{{}^{\prime}}=(1-\gamma)\tau_{(n+1)}$ Denoise
     ${\bm{x}}_{\tau_{n}\rightarrow\tau_{n}^{{}^{\prime}}}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}(\tilde{{\bm{x}}}_{\tau_{n}},\tau_{n},\tau_{n}^{{}^{\prime}})$ Diffuse
     $\tilde{{\bm{x}}}_{\tau_{(n+1)}}\leftarrow\frac{\alpha_{\tau_{(n+1)}}}{\alpha_{\tau_{n}^{{}^{\prime}}}}{\bm{x}}_{\tau_{n}\rightarrow\tau_{n}^{{}^{\prime}}}+\sqrt{1-\frac{\alpha_{\tau_{(n+1)}}^{2}}{\alpha_{\tau_{n}^{{}^{\prime}}}^{2}}}{\bm{z}}$
  end for
  ${\bm{x}}_{\tau_{N}\rightarrow 0}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}(\tilde{{\bm{x}}}_{\tau_{N}},\tau_{N},0)$
  Output: ${\bm{x}}_{\tau_{N}\rightarrow 0}$
Appendix C Implementation Details
C.1 Dataset
We trained TCD on the LAION5B High-Res dataset (Schuhmann et al., 2022). Images with an aesthetic score (Schuhmann, 2022) lower than 5.8 were filtered out. For training, all images were initially resized to 1024 pixels by the shorter side and subsequently randomly cropped to dimensions of $1024\times 1024$.
C.2 Hyper-parameters
In our experiments, we utilized the AdamW optimizer with $\beta_{1}=0.9,\beta_{2}=0.999$, and a weight decay of 0.01. For TCF(2) parameterization, the itermidiate timestep $u$ in Equation 18 was set as $u:=t_{\lambda}(\lambda_{s}+\lambda_{t})$, where $t_{\lambda}$ is the inverse function of $\lambda_{t}$. For TCF(S+) parameterization, to encode the ending time $s$ into the TCD model, we create sinusoidal timestep embeddings matching the implementation in (Ho et al., 2020). Subsequently, we project it with a simple MLP, integrating it into the TCD backbone by adding the projected $s$-embedding to the start time $t$-embedding, following the approach in the original Stable Diffusion models (Rombach et al., 2022). We choose the DDIM (Song et al., 2020a) as the ODE solver and set the skipping step $k=20$ in Equation 21. During training, we initialize the trajectory consistency function with the same parameters as the teacher diffusion model. The guidance scale for guided distillation (Meng et al., 2023) was set to $[2,14]$. We did not utilize the EDM’s skip scale and output scale, as we found it did not provide any benefit. For TCF(1) and TCF(2), the batch size was set to 256, and the learning rate was set to 4.5e-6, and the LoRA rank was 64. We did not use EMA for TCF(1) and TCF(2). For TCF(S+), the batch size was set to 192, the learning rate was set to 4.0e-6. We did not use LoRA for TCF(S+) but directly fine-tuned the U-Net, and the EMA decay rate was set to 0.95.
C.3 Hardware and Efficiency
All our models were trained using 8 80G A800 units. TCF(1) was trained for 3000 iterations, taking 15 hours, demonstrating high training efficiency. TCF(2) employed the same parameterization method as TCF(1) during training. TCF(S+) was trained for 43,000 iterations, spanning 5 days and 20 hours. The low convergence rate of TCF(S+) may be attributed to the introduction of additional parameters, resulting in low distillation efficiency of teacher information. We defer the exploration of improved design for enhancing the efficiency of TCF(S+) to future work. The inference time in Figure 1 was tested on an A800, averaged over 16 prompts.
Appendix D Proofs
D.1 Proof of Corollary 3.1
Before we prove the corollary 3.1, we assume the following mild conditions on the data distribution $p_{\rm{data}}$ (Lee et al., 2023; Chen et al., 2022):
Assumption D.1.
The data distribution has finite second moment, that is, $\mathbb{E}_{{\bm{x}}_{0}\sim p_{\rm{data}}}[\|{\bm{x}}_{0}\|_{2}^{2}]=m_{2}^{2}<\infty$.
Assumption D.2.
The score function $\nabla\log p_{t}({\bm{x}})$ is Lipschitz on the variable ${\bm{x}}$ with Lipschitz constant $L_{s}\geq 1$, $\forall t\in[0,T]$.
We further assume bounds on the score estimation error and consistency error (Song et al., 2023; Lyu et al., 2023):
Assumption D.3.
Assume $\mathbb{E}_{{\bm{x}}_{t_{n}}\sim p_{t_{n}}}[\|{\bm{s}}_{\bm{\phi}}({\bm{x}}_{t_{n}},t_{n})-\nabla\log p_{t_{n}}({\bm{x}}_{t_{n}})\|_{2}^{2}]\leq\varepsilon_{\text{sc}}^{2},\forall n\in[\![1,N]\!]$.
Assumption D.4.
Assume $\mathbb{E}_{{\bm{x}}_{t_{n}}\sim p_{t_{n}}}[\|{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n+1}},t_{n+1})-{\bm{f}}_{{\bm{\theta}}}(\hat{\bm{x}}_{t_{n}}^{\bm{\phi}},t_{n})\|_{2}^{2}]\leq\varepsilon_{\text{cm}}^{2}(t_{n+1}-t_{n})^{2},\forall n\in[\![1,N-1]\!]$
Assumption D.5.
The consistency model ${\bm{f}}_{\bm{\theta}}({\bm{x}},t_{n})$ is Lipschitz on the variable ${\bm{x}}$ with Lipschitz constant $L_{f}>1,~{}\forall n\in[\![1,N]\!].$
We also follow the (Lyu et al., 2023) to assume a discrete schedule:
Assumption D.6.
Assume the discretisation schedule $0<\delta=t_{1}<t_{2}<\cdots<t_{N}=T$, $h_{k}=t_{k+1}-t_{k}$ to Equation 3 is divided into two stages:
  1. 1.
    $h_{k}\equiv h$ for all $k\in[\![N_{1},N-1]\!],$ and $(N-N_{1}-1)h<T\leq(N-N_{1})h$;
  2. 2.
    $h_{k}=2^{-(N_{1}-k)}h=\frac{h_{k+1}}{2}$ for $k\in[\![1,N_{1}-1]\!]$, $N_{1}$ satisfies $h_{2}=2^{-(N_{1}-2)}h\leq 2\delta.$
Then we introduce some properties and a lemma borrowed from (Lyu et al., 2023), which shows that TV error can be bounded after a small time OU regularisation.
The forward OU process can be defined as follows:
$\mathrm{d}{\bm{x}}_{t}=-{\bm{x}}_{t}\mathrm{d}t+\sqrt{2}\mathrm{d}{\bm{w}}_{t},$ (26)
and denote the Markov kernel $P_{\text{OU}}^{s}$ to be defined by Equation 26, that is, if ${\bm{x}}_{t}\sim p$ for some $p$ be a distribution over $\mathbb{R}^{d}$, then ${\bm{x}}_{t+s}\sim pP_{\text{OU}}^{s}$.
Under Assumptions D.1-D.6, when $T>L_{s}^{-1}$, the one-step generating error is bounded as follows:
$W_{2}({\bm{f}}_{{\bm{\theta}},T}\sharp\mathcal{N}({\bm{0}},{\bm{I}}_{d}),p_{\rm{data}})\lesssim(d^{\frac{1}{2}}\vee m_{2})L_{f}e^{-T}+T(\varepsilon_{\text{cm}}+L_{f}\varepsilon_{\text{sc}}+L_{f}L_{s}^{\frac{3}{2}}d^{\frac{1}{2}}h)+(d^{\frac{1}{2}}\vee m_{2})\delta^{\frac{1}{2}}.$ (27)
Lemma D.7.
For any two distributions $p$ and $q$, running the OU process in Equation 26 for $p,q$ individually with time $\tau>0$, the following TV distance bound holds,
[…]
Proof.
Denote $\psi_{\sigma^{2}}({\bm{y}})$ as the density function to the normal distribution $\mathcal{N}({\bm{0}},\sigma^{2}{\bm{I}}_{d})$. We write the $\text{TV}(pP_{\text{OU}}^{\tau},qP_{\text{OU}}^{\tau})$ into integral form as:
$\text{TV}(pP_{\text{OU}}^{\tau},qP_{\text{OU}}^{\tau})$ $=\frac{1}{2}\int_{\mathbb{R}^{d}}|(pP_{\text{OU}}^{\tau})({\bm{x}})-(qP_{\text{OU}}^{\tau})({\bm{x}})|\mathrm{d}{\bm{x}}$
$=\frac{1}{2}\int_{\mathbb{R}^{d}}\left|\int_{\mathbb{R}^{d}}p({\bm{y}})\psi_{1-e^{-2\tau}}({\bm{x}}-e^{-\tau}{\bm{y}})\mathrm{d}{\bm{y}}-\int_{\mathbb{R}^{d}}q({\bm{z}})\psi_{1-e^{-2\tau}}({\bm{x}}-e^{-\tau}{\bm{z}})\mathrm{d}{\bm{z}}\right|\mathrm{d}{\bm{x}}.$ (28)
Take a coupling $\gamma\in\Gamma(p,q)$, then
$\int_{\mathbb{R}^{d}}\gamma({\bm{y}},{\bm{z}})\mathrm{d}{\bm{z}}=p({\bm{y}}),$
$\int_{\mathbb{R}^{d}}\gamma({\bm{y}},{\bm{z}})\mathrm{d}{\bm{y}}=q({\bm{z}}),$ (29)
we have
$\text{TV}(pP_{\text{OU}}^{\tau},qP_{\text{OU}}^{\tau})$ $=\frac{1}{2}\int_{\mathbb{R}^{d}}\left|\int_{\mathbb{R}^{\mathrm{d}\times d}}\gamma({\bm{y}},{\bm{z}})[\psi_{1-e^{-2\tau}}({\bm{x}}-e^{-\tau}{\bm{y}})-\psi_{1-e^{-2\tau}}({\bm{x}}-e^{-\tau}{\bm{z}})]\mathrm{d}{\bm{y}}\mathrm{d}{\bm{z}}\right|\mathrm{d}{\bm{x}}.$
$\leq\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d\times d}}\gamma({\bm{y}},{\bm{z}})\left|\psi_{1-e^{-2\tau}}({\bm{x}}-e^{-\tau}{\bm{y}})-\psi_{1-e^{-2\tau}}({\bm{x}}-e^{-\tau}{\bm{z}})\right|\mathrm{d}{\bm{y}}\mathrm{d}{\bm{z}}\mathrm{d}{\bm{x}}$
$=\int_{\mathbb{R}^{d\times d}}\gamma({\bm{y}},{\bm{z}})\left(\frac{1}{2}\int_{\mathbb{R}^{d}}\left|\psi_{1-e^{-2\tau}}({\bm{x}}-e^{-\tau}{\bm{y}})-\psi_{1-e^{-2\tau}}({\bm{x}}-e^{-\tau}{\bm{z}})\right|\mathrm{d}{\bm{x}}\right)\mathrm{d}{\bm{y}}\mathrm{d}{\bm{z}}$
$=\int_{\mathbb{R}^{d\times d}}\gamma({\bm{y}},{\bm{z}})\text{TV}(\psi_{1-e^{-2\tau}}(\cdot-e^{-\tau}{\bm{y}}),\psi_{1-e^{-2\tau}}(\cdot-e^{-\tau}{\bm{z}}))\mathrm{d}{\bm{y}}\mathrm{d}{\bm{z}}$
$\leq\int_{\mathbb{R}^{d\times d}}\gamma({\bm{y}},{\bm{z}})\sqrt{\frac{1}{2}\text{KL}(\psi_{1-e^{-2\tau}}(\cdot-e^{-\tau}{\bm{y}})\|\psi_{1-e^{-2\tau}}(\cdot-e^{-\tau}{\bm{z}}))}\mathrm{d}{\bm{y}}\mathrm{d}{\bm{z}}$
$=\int_{\mathbb{R}^{d\times d}}\gamma({\bm{y}},{\bm{z}})\frac{1}{2}\sqrt{\frac{e^{-2\tau}}{1-e^{-2\tau}}\|{\bm{y}}-{\bm{z}}\|_{2}^{2}}\mathrm{d}{\bm{y}}\mathrm{d}{\bm{z}}$
$=\frac{1}{2\sqrt{e^{2\tau}-1}}\int_{\mathbb{R}^{d\times d}}\gamma({\bm{y}},{\bm{z}})\|{\bm{y}}-{\bm{z}}\|_{2}\mathrm{d}{\bm{y}}\mathrm{d}{\bm{z}}$ (30)
Noting $\frac{1}{2\sqrt{e^{2\tau}-1}}\leq\frac{1}{2\sqrt{2\tau}}$, and taking $\gamma$ over all coupling $\Gamma(p,q)$, we have
$\text{TV}(pP_{\text{OU}}^{\tau},qP_{\text{OU}}^{\tau})\lesssim\frac{1}{\sqrt{\tau}}W_{1}(p,q)\leq\frac{1}{\sqrt{\tau}}W_{2}(p,q)$
Let $q$ be the output of consistency models, either the one step consistency sampling result or the $k$-th multistep consistency sampling result. To control the TV error, we smooth the generated sample by the forward OU process with a small time that is the same as the early stopping time $\delta$, and then we can get the TV distance between $q_{k}P_{\text{OU}}^{\delta}$ and $p_{\rm{data}}$.
Proof of Corollary 3.1.
According to the triangular inequality, Lemma D.7 and Equation 27,
$\text{TV}(q_{1}P_{\text{OU}}^{\delta},p_{\rm{data}})$ $\leq\text{TV}(q_{1}P_{\text{OU}}^{\delta},p_{\delta}P_{\text{OU}}^{\delta})+\text{TV}(p_{\delta}P_{\text{OU}}^{\delta},p_{\rm{data}})$
$\lesssim\frac{1}{\sqrt{\delta}}W_{2}(q_{1},p_{\delta})+\text{TV}(p_{2\delta},p_{\rm{data}})$
$\lesssim\frac{1}{\sqrt{\delta}}\left((d^{\frac{1}{2}}\vee m_{2})L_{f}e^{-T}+T(\varepsilon_{\text{cm}}+L_{f}\varepsilon_{\text{sc}}+L_{f}L_{s}^{\frac{3}{2}}d^{\frac{1}{2}}h)\right)+\text{TV}(p_{2\delta},p_{\rm{data}}).$ (31)
Note that if we take $\delta\asymp\frac{\varepsilon^{2}}{L_{s}^{2}(d\vee m_{2}^{2})}$, then by Lemma 6.4, (Lee et al., 2023), $\text{TV}(p_{2\delta},p_{\rm{data}})\leq\varepsilon$, this concludes that
$\text{TV}(q_{1}P_{\text{OU}}^{\delta},p_{\rm{data}})\lesssim\frac{L_{s}(d^{\frac{1}{2}}\vee m_{2})}{\varepsilon}[(\log L_{f}+\frac{T}{2^{k}})(\varepsilon_{\text{cm}}+L_{f}\varepsilon_{\text{sc}}+L_{f}L_{s}^{\frac{3}{2}}d^{\frac{1}{2}}h)+\frac{(d^{\frac{1}{2}}\vee m_{2})L_{f}}{2^{k}e^{T}}]+\varepsilon.$
Similarly for $q_{k}$, if we take $\delta\asymp\frac{\varepsilon^{2}}{L_{s}^{2}(d\vee m_{2}^{2})}$
$\text{TV}(q_{k}P_{\text{OU}}^{\delta},p_{\rm{data}})$ (32)
$\leq$ $\text{TV}(q_{k}P_{\text{OU}}^{\delta},p_{\delta}P_{\text{OU}}^{\delta})+\text{TV}(p_{\delta}P_{\text{OU}}^{\delta},p_{\rm{data}})$
$\lesssim$ $\frac{L_{s}(d^{\frac{1}{2}}\vee m_{2})}{\varepsilon}[(\log L_{f}+\frac{T}{2^{k}})(\varepsilon_{\text{cm}}+L_{f}\varepsilon_{\text{sc}}+L_{f}L_{s}^{\frac{3}{2}}d^{\frac{1}{2}}h)+\frac{(d^{\frac{1}{2}}\vee m_{2})L_{f}}{2^{k}e^{T}}]+\varepsilon.$ (33)
D.2 Proof of Theorem 4.1
The proof closely follows Song et al. (2023) and relies on induction, drawing parallels with the classic proof of global error bounds for numerical ODE solvers (Prasad, 1990).
Proof.
For simplicity of notation, we denote $\bm{\theta^{*}}$ as $\bm{\theta}$. From $\mathcal{L}_{\rm{TCD}}^{N}(\bm{\theta},\bm{\theta};\bm{\phi})=0$, we have
$\mathcal{L}^{N}_{\rm{TCD}}=\mathbb{E}\left[\omega(t_{n},t_{m})\left\|{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n+1}},t_{n+1},t_{m})-{\bm{f}}_{\bm{\theta}}(\hat{{\bm{x}}}_{t_{n}}^{\bm{\phi}},t_{n},t_{m})\right\|^{2}_{2}\right]=0.$ (34)
According to the definition, it follows that $p_{t_{n}}({\bm{x}}_{t_{n}})>0$ for every $x_{t_{n}}$ and $1\leqslant n\leqslant N$. Therefore,  Equation 34 entails
$\omega(t_{n},t_{m})\left\|{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n+1}},t_{n+1},t_{m})-{\bm{f}}_{\bm{\theta}}(\hat{{\bm{x}}}_{t_{n}}^{\bm{\phi}},t_{n},t_{m})\right\|^{2}_{2}\equiv 0.$ (35)
Because $\lambda(\cdot)>0$ and $\left\|{\bm{x}},{\mathbf{y}}\right\|^{2}_{2}=0\Leftrightarrow{\bm{x}}={\mathbf{y}}$, this further implies that
${\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n+1}},t_{n+1},t_{m})\equiv{\bm{f}}_{\bm{\theta}}(\hat{{\bm{x}}}_{t_{n}}^{\bm{\phi}},t_{n},t_{m}).$ (36)
Let ${\bm{e}}_{n}^{s}$ represent the error vector that arises when predicting $x_{s}$ at $t_{n}$, which is defined as
${\bm{e}}_{n,m}:={\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n}},t_{n},t_{m})-{\bm{f}}({\bm{x}}_{t_{n}},t_{n},t_{m};\bm{\phi}).$
We can easily derive the following recursion relation
${\bm{e}}_{n+1,m}$ $={\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n+1}},t_{n+1},t_{m})-{\bm{f}}({\bm{x}}_{t_{n+1}},t_{n+1},t_{m};\bm{\phi})$ (37)
$={\bm{f}}_{\bm{\theta}}(\hat{{\bm{x}}}_{t_{n}}^{\bm{\phi}},t_{n},t_{m})-{\bm{f}}({\bm{x}}_{t_{n}},t_{n},t_{m};\bm{\phi})$
$={\bm{f}}_{\bm{\theta}}(\hat{{\bm{x}}}_{t_{n}}^{\bm{\phi}},t_{n},t_{m})-{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n}},t_{n},t_{m})+{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n}},t_{n},t_{m})-{\bm{f}}({\bm{x}}_{t_{n}},t_{n},t_{m};\bm{\phi})$
$={\bm{f}}_{\bm{\theta}}(\hat{{\bm{x}}}_{t_{n}}^{\bm{\phi}},t_{n},t_{m})-{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n}},t_{n},t_{m})+{\bm{e}}_{n,m}.$
Because ${\bm{f}}_{\theta}(\cdot,t_{n},t_{m})$ has Lipschitz constant $L$, we have
$\lVert{\bm{e}}_{n+1,m}\rVert_{2}$ $\leqslant\lVert{\bm{e}}_{n,m}\rVert_{2}+L\left\lVert\hat{{\bm{x}}}_{t_{n}}^{\bm{\phi}}-{\bm{x}}_{t_{n}}\right\rVert_{2}$
$=\lVert{\bm{e}}_{n,m}\rVert_{2}+L\cdot\mathcal{O}((t_{n+1}-t_{n})^{p+1})$
$=\lVert{\bm{e}}_{n,m}\rVert_{2}+\mathcal{O}((t_{n+1}-t_{n})^{p+1}.$
In addition, we observe that ${\bm{e}}_{m,m}=\bm{0}$, because
${\bm{e}}_{m,m}$ $={\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{m}},t_{m},t_{m})-{\bm{f}}({\bm{x}}_{t_{m}},t_{m},t_{m};\bm{\phi})$
=$\overset{(i)}{=}{\bm{x}}_{t_{m}}-{\bm{f}}({\bm{x}}_{t_{m}},t_{m},t_{m};\bm{\phi})$
=$\overset{(i)}{=}{\bm{x}}_{t_{m}}-{\bm{x}}_{t_{m}}$
$=\bm{0},$
here $(i)$ is true because trajectory consistency distillation restricts the boundary condition for the parameterized model such that ${\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{m}},t_{m},t_{m})={\bm{x}}_{t_{m}}$, and $(ii)$ is entailed by the definition of trajectory function ${\bm{f}}(\cdot,\cdot,\cdot;\bm{\phi})$ in  Equation 13. This allows us to perform induction on the recursion formula Equation 37 to obtain
$\lVert{\bm{e}}_{n,m}\rVert_{2}$ $\leqslant\lVert{\bm{e}}_{m,m}\rVert_{2}+\sum_{k=m}^{n-1}\mathcal{O}((t_{k+1}-t_{k})^{p+1}$
$=\sum_{k=m}^{n-1}\mathcal{O}((t_{k+1}-t_{k})^{p+1}$
$=\sum_{k=m}^{n-1}(t_{k+1}-t_{k})\mathcal{O}((t_{k+1}-t_{k})^{p}$
$\leqslant\sum_{k=m}^{n-1}(t_{k+1}-t_{k})\mathcal{O}((\Delta t)^{p}$
$=\mathcal{O}((\Delta t)^{p}\sum_{k=m}^{n-1}(t_{k+1}-t_{k})$
$=\mathcal{O}((\Delta t)^{p}(t_{n}-t_{m}),$
which completes the proof. ∎
D.3 Proof of  Theorem 4.2
In this section, our derivation mainly borrows the proof from (Kim et al., 2023; Chen et al., 2022).
For DDPM (Chen et al., 2022),
Theorem D.8.
Suppose that Assumptions D.2D.1, and D.3 hold. Let $q_{T}$ be the output of the DDPM algorithm at time $T$, and suppose that the step size $h:=T/N$ satisfies $h\lesssim 1/L_{s}$, where $L_{s}\geq 1$ is the Lipschitz constant of score function. Then, it holds that
[…]
For simplicity, we assume the optimal TCD, $f_{\bm{\theta}^{*}}\equiv f$ with a well-learned $\bm{\theta}^{*}$, which recovers the true $f(\cdot,t,s)$ function. We establish that the density propagated by this optimal TCD model from any time $t$ to a subsequent time $s$ aligns with the predefined density determined by the fixed forward process.
We now present the proposition ensuring the alignment of the transited density.
Proposition D.9.
[…]
  1. 1.
    Linear growth: $\|\nabla\log p_{t}(\mathbf{x})\|_{2}\leq L(t)(1+\|\mathbf{x}\|_{2})$, for all $\mathbf{x}\in\mathbb{R}^{D}$
  2. 2.
    Lipschitz: $\|\nabla\log p_{t}(\mathbf{x})-\nabla\log p_{t}(\mathbf{y})\|_{2}\leq L(t)\|\mathbf{x}-\mathbf{y}\|_{2}$, for all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{D}$.
Then for any $t\in[0,T]$ and $s\in[0,t]$, $p_{s}=\mathcal{T}_{t\rightarrow s}\sharp p_{t}$.
This theorem guarantees that by learning the optimal TCD, which possesses complete trajectory information, we can retrieve all true densities at any time using TCD.
Define $\mathcal{T}_{t\rightarrow s}$ as the oracle transition mapping from $t$ to $s$ via the diffusion process in Equation 1. Let $\mathcal{T}_{t\rightarrow s}^{\bm{\theta}^{*}}(\cdot)$ represent the transition mapping from the optimal TCD model, and $\mathcal{T}_{t\rightarrow s}^{\bm{\phi}}(\cdot)$ represent the transition mapping from the empirical probability flow ODE. Since all processes start at point $T$ with an initial probability distribution $p_{T}$ and $\mathcal{T}_{t\rightarrow s}^{\bm{\theta}^{*}}(\cdot)=\mathcal{T}_{t\rightarrow s}^{\bm{\phi}}(\cdot)$Theorem D.8 and $\mathcal{T}_{T\rightarrow t}\sharp p_{T}=p_{t}$ from Proposition D.9 tell us that for $t>s$
$\text{TV}\left(\mathcal{T}_{t\rightarrow s}\sharp p_{t},\mathcal{T}_{t\rightarrow s}^{\bm{\theta}^{*}}\sharp p_{t}\right)=\text{TV}\left(\mathcal{T}_{t\rightarrow s}\sharp p_{t},\mathcal{T}_{t\rightarrow s}^{\bm{\phi}}\sharp p_{t}\right)=\mathcal{O}(\sqrt{t-s}).$ (38)
$\text{TV}\left(\mathcal{T}_{t\rightarrow 0}\mathcal{T}_{(1-\gamma)t\rightarrow t}\mathcal{T}_{T\rightarrow(1-\gamma)t}\sharp p_{T},\mathcal{T}_{t\rightarrow 0}^{\bm{\theta}^{*}}\mathcal{T}_{(1-\gamma)t\rightarrow t}\mathcal{T}_{T\rightarrow(1-\gamma)t}^{\bm{\theta}^{*}}\sharp p_{T}\right)$
$\overset{(a)}{\leq}$ $\text{TV}\left(\mathcal{T}_{t\rightarrow 0}\mathcal{T}_{(1-\gamma)t\rightarrow t}\mathcal{T}_{T\rightarrow(1-\gamma)t}\sharp p_{T},\mathcal{T}_{t\rightarrow 0}^{\bm{\theta}^{*}}\mathcal{T}_{(1-\gamma)t\rightarrow t}\mathcal{T}_{T\rightarrow(1-\gamma)t}\sharp p_{T}\right)$
$+$ $\text{TV}\left(\mathcal{T}_{t\rightarrow 0}^{\bm{\theta}^{*}}\mathcal{T}_{(1-\gamma)t\rightarrow t}\mathcal{T}_{T\rightarrow(1-\gamma)t}\sharp p_{T},\mathcal{T}_{t\rightarrow 0}^{\bm{\theta}^{*}}\mathcal{T}_{(1-\gamma)t\rightarrow t}\mathcal{T}_{T\rightarrow(1-\gamma)t}^{\bm{\theta}^{*}}\sharp p_{T}\right)$
=$\overset{(b)}{=}$ $\text{TV}\left(\mathcal{T}_{t\rightarrow 0}\mathcal{T}_{T\rightarrow t}\sharp p_{T},\mathcal{T}_{t\rightarrow 0}^{\bm{\theta}^{*}}\mathcal{T}_{T\rightarrow t}\sharp p_{T}\right)+\text{TV}\left(\mathcal{T}_{T\rightarrow(1-\gamma)t}\sharp p_{T},\mathcal{T}_{T\rightarrow(1-\gamma)t}^{\bm{\theta}^{*}}\sharp p_{T}\right)$
=$\overset{(c)}{=}$ $\text{TV}\left(\mathcal{T}_{t\rightarrow 0}\sharp p_{t},\mathcal{T}_{t\rightarrow 0}^{\bm{\theta}^{*}}\sharp p_{t}\right)+\text{TV}\left(\mathcal{T}_{T\rightarrow(1-\gamma)t}\sharp p_{T},\mathcal{T}_{T\rightarrow(1-\gamma)t}^{\bm{\theta}^{*}}\sharp p_{T}\right)$
=$\overset{(d)}{=}$ $\mathcal{O}(\sqrt{t})+\mathcal{O}(\sqrt{T-(1-\gamma)t}).$
Here (a) is obtained from the triangular inequality, (b) and (c) are due to $\mathcal{T}_{(1-\gamma)t\rightarrow t}\mathcal{T}_{T\rightarrow(1-\gamma)t}=\mathcal{T}_{T\rightarrow t}$ and $\mathcal{T}_{T\rightarrow t}\sharp p_{T}=p_{t}$ from Proposition D.9, and (d) comes from Equation 38.
Appendix E Additional Results
E.1 More Samples with Different NFEs
In Figure 8, we present additional samples synthesized by LCM and TCD with varying NFEs. These samples consistently demonstrate that the details of images generated by LCM tend to vanish as the NFEs increase, whereas TCD either maintains or enhances image details at higher NFEs.
E.2 More Samples with Different Stochastic Parameter $\gamma$
In Figure 9, we showcase additional samples generated by TCD with varying stochastic parameter $\gamma$, demonstrating that the visual complexity and quality of the image gradually improve as $\gamma$ increases.
E.3 More Comparisons Results
In Figure 10 to Figure 14, we present additional examples for comparison with state-of-the-art methods. These samples illustrate that our TCD can generate high-quality images with few NFEs, outperforming LCM. Additionally, it can produce more detailed and high-definition images with sufficient NFEs, even surpassing the performance of DPM-Solver++(2S).
E.4 More samples from TCD’s Versatility Testing
To ascertain the versatility of TCD, we test it on a wide range of models, including the popular community model Animagine XL V3  [^6] The checkpoint of Animagine can be found in https://civitai.com/models/260267/animagine-xl-v3, LoRA Papercut XL  [^7] The checkpoint of Papercut can be found in https://civitai.com/models/122567/papercut-sdxl, Depth ControlNet  [^8] Depth ControlNet: https://huggingface.co/diffusers/controlnet-depth-sdxl-1.0, Canny ControlNet  [^9] Canny ControlNet: https://huggingface.co/diffusers/controlnet-canny-sdxl-1.0, and IP-Adapter  [^10] IP-Adapter: https://github.com/tencent-ailab/IP-Adapter. In Figure 15 to Figure 19, we observe that TCD can be directly applied to all these models, accelerating their sampling with only 4 steps. It is worth noting that, in this experiment, all models share the same TCD LoRA parameters.
Figure 8: More samples with different NFEs.
Figure 9: More samples with different stochastic parameter $\gamma$.
Figure 10: More comparisons results.
Figure 11: More comparisons results.
Figure 12: More comparisons results.
Figure 13: More comparisons results.
Figure 14: More comparisons results.
Figure 15: Qualitative results of TCD using different base models: SDXL and Animagine XL V3. It is worth noting that we employed the same TCD parameters for both models. All samples were generated using 4 steps. In each subfigure, the top row corresponds to TCD + SDXL, and the bottom row corresponds to TCD + Animagine XL V3.
Figure 16: Qualitative results of TCD with and without Papercut XL LoRA. We used the same TCD parameters. All samples are generated using 4 steps. In each subfigure, the top row corresponds to TCD without Papercut LoRA, and the bottom row corresponds to TCD with Papercut LoRA. The Lora scale of Papercut was set to 1.0 in the experiments.
Figure 17: Qualitative results of TCD with Depth ControlNet. All samples are generated using 4 steps.
Figure 18: Qualitative results of TCD with Canny ControlNet. All samples are generated using 4 steps.
Figure 19: Qualitative results of TCD with IP-Adapter. All samples are generated using 4 steps.

### v2

@@
-Trajectory Consistency Distillation
+Trajectory Consistency Distillation: Improved Latent Consistency Distillation by Semi-Linear Consistency Function with Trajectory Mapping
 Abstract
@@
 However, we observed that LCM struggles to generate images with both clarity and detailed intricacy.
-To address this limitation, we initially delve into and elucidate the underlying causes.
-Our investigation identifies that the primary issue stems from errors in three distinct areas.
 Consequently, we introduce Trajectory Consistency Distillation (TCD), which encompasses trajectory consistency function and strategic stochastic sampling.
-The trajectory consistency function diminishes the distillation errors by broadening the scope of the self-consistency boundary condition and endowing the TCD with the ability to accurately trace the entire trajectory of the Probability Flow ODE.
-Additionally, strategic stochastic sampling is specifically designed to circumvent the accumulated errors inherent in multi-step consistency sampling, which is meticulously tailored to complement the TCD model.
+The trajectory consistency function diminishes the parameterisation and distillation errors by broadening the scope of the self-consistency boundary condition with trajectory mapping and endowing the TCD with the ability to accurately trace the entire trajectory of the Probability Flow ODE in semi-linear form with an Exponential Integrator.
+Additionally, strategic stochastic sampling provides explicit control of stochastic and circumvents the accumulated errors inherent in multi-step consistency sampling.
 Experiments demonstrate that TCD not only significantly enhances image quality at low NFEs but also yields more detailed results compared to the teacher model at high NFEs.
@@
 It is noteworthy that all these Consistency-Type Models still allow for striking a balance between computation and sample quality using Multistep Consistency Sampling (Song et al., 2023).
-In particular, allocating additional compute for more iterations can theoretically yield samples of higher quality.
+There are also some works that focus on changing the interval of the consistency function or the distillation learning protocol (Gu et al., 2023; Berthelot et al., 2023; Heek et al., 2024), leading to enhanced outcomes.
 Figure 1: Comparison between TCD and other state-of-the-art methods.
@@
 Additionally, Li et al.
-(2023) showed evidence that during the process of multi-step sampling, discrete errors accumulate over iterations, ultimately causing the generated images to deviate from the target distribution.
-Drawing inspiration from the observation, we first meticulously examine the training procedure and delve into the multistep consistency sampling procedure to figure out the root cause.
-Our investigation revealed that the issue stems from the cumulative errors inherent in multi-step sampling processes.
-These errors predominantly originate from three sources: 1) the estimation errors in the original score matching model, 2) the distillation errors in the consistency model, and 3) the discretisation errors accumulated during the sampling phase.
-Collectively, these errors significantly undermine the efficacy of the multistep sampling consistency model, resulting in a performance that falls substantially short of expectations.
-To suppress these errors, we introduce Trajectory Consistency Distillation (TCD), as summarized in Figure 2, which comprises two key elements: trajectory consistency function and strategic stochastic sampling.
-Specifically, inspired by the form of exponential integrators, the trajectory consistency function (TCF) diminishes distillation errors by expanding the boundary conditions of the consistency model and enabling seamless transitions at any point along the trajectory governed by the PF ODE, as demonstrated in Figure 1(a).
-Furthermore, strategic stochastic sampling (SSS) suppresses the accumulated discretisation error and estimation error according to the narrowed bi-directional iterations as shown in Figure 1(b).
+(2023) showed evidence that during the process of multi-step sampling, discretisation errors accumulate over iterations, ultimately causing the generated images to deviate from the target distribution.
+Recently, in order to balance the inference steps and the sample quality during inference, CTM (Kim et al., 2023) proposed a unified framework that takes advantage of the sampling methods of Distillation Models (Salimans & Ho, 2022; Song et al., 2023) and SGMs according to a novel representation of denoisers and analyzed the errors during sampling.
+In this paper, we focus on improving the generation quality of LCMs, especially reducing the errors arising during parameterisation process of LCMs.
+We introduce Trajectory Consistency Distillation (TCD), as summarized in Figure 2, which comprises two key elements: trajectory consistency function and strategic stochastic sampling.
+Specifically, inspired by the form of exponential integrators (Lu et al., 2022a, b), the trajectory consistency function (TCF) diminishes parameterisation and distillation errors by expanding the boundary conditions of the consistency model and enabling seamless transitions at any point along the trajectory governed by the PF ODE, as demonstrated in Figure 2(a).
+Furthermore, strategic stochastic sampling (SSS) allows control of stochasticity and suppresses the accumulated discretisation error and estimation error according to the narrowed bi-directional iterations as shown in Figure 2(b).
 Experiments show that TCD can significantly enhance the quality of images generated by LCM, surpassing it in performance.
@@
 SDXL with DPMSolver++) when sampling with sufficient iterations (e.g., 20 NFEs).
+2 Related Works
+Diffusion Models.
+Diffusion Models (DMs), also known as Score-based generative models, have shown superior performance in various generative fields.
+Chen et al.
+(2022) provides theoretical convergence guarantees, implying that DMs can efficiently sample from essentially any realistic data distribution under minimal data assumptions.
+ADM (Dhariwal & Nichol, 2021) firstly shows the potential of DMs to outperform GANs.
+EDM (Karras et al., 2022) further elucidates the design space of DMs, clearly separating the concrete design choices, deriving best practices for the sampling process, and improving the training dynamics, thereby drastically improving the results.
+Text-conditional Diffusion Models.
+DMs have specially achieved great success in the area of text-to-image synthesis (Nichol et al., 2022; Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022; Balaji et al., 2022; Podell et al., 2023).
+To reduce computational cost, diffusion models typically operate within the latent space (Rombach et al., 2022; Podell et al., 2023) or include separate super-resolution steps (Ramesh et al., 2022; Saharia et al., 2022; Balaji et al., 2022).
+The integration of classifier-free diffusion guidance (Ho & Salimans, 2022; Dieleman, 2022) during the sampling process dramatically improves samples produced by conditional diffusion models at almost no additional cost.
+Fast Sampling of DMs.
+DMs exhibit great generative capabilities but are bottlenecked by their slow sampling speed.
+Various numerical methods driven by the SDE mathematical underpinnings underlying DMs have been proposed for accelerating DM sampling.
+DDIM (Song et al., 2020a) has originally shown promise for few-step sampling.
+Other works involve predictor-corrector samplers (Song et al., 2020b; Karras et al., 2022), exponential integrators (Lu et al., 2022a; Zhang & Chen, 2022; Lu et al., 2022b), and automated methods to adjust solvers (Kong & Ping, 2021; Jolicoeur-Martineau et al., 2021).
+Among them, the semi-linear structure utilizing exponential integrators (Lu et al., 2022a; Zhang & Chen, 2022; Lu et al., 2022b) can derive accurate solutions for the PF-ODEs.
+By separating the linear and nonlinear parts, it skillfully bypasses the inherent discretization errors in the linear items, thereby obtaining exact analytical solutions for the linear term.
+Another series of research works exemplified by distillation techniques (Luhman & Luhman, 2021; Salimans & Ho, 2022; Meng et al., 2023).
+These methods distill knowledge from pretrained models into few-step samplers, representing an efficient solution for few NFEs.
+However, they may experience a lengthy and costly process that requires huge amounts of data and suffer from slow convergence.
+Consistency Models.
+To overcome the limitations of current fast samplers, Song et al.
+(2023) proposes Consistency Models (CMs) learning a direct mapping from noise to data built on the top of the trajectory of PF ODE, achieving one-step generation while allowing multi-step sampling to trade compute for quality.
+Lyu et al.
+(2023) provide the first convergence guarantees for CMs under moderate assumptions.
+Kim et al.
+(2023) proposes a universal framework for CMs and DMs samplings with an “anytime to anytime” decoder, which is able to predict both infinitesimally small steps and arbitrary long step jumps.
+Together with a $\gamma$-sampler, it achieves a balance between sampling quality and computational resources.
+Both CTM and TCD eliminate constraints on predictions at the origin during distillation, but from distinct viewpoints, leading to different solutions.
+Furthermore, CTM provides $\gamma$ sampling with a well-founded theoretical basis in the view of multistep sampling, while our experiments disseminate its theoretical findings.
+Further comments about CTM and TCD can be found in Appendix A.
+Latent Consistency Models (LCMs) (Luo et al., 2023a) integrate consistency distillation with latent diffusion models (Rombach et al., 2022) to achieve impressive performance in accelerating text-to-image synthesis.
+LCM-LoRA (Luo et al., 2023b) further improves the training efficiency and versatility of LCMs by introducing LoRA (Hu et al., 2021) into the distillation process.
+Several studies have concentrated on modifying the interval of the consistency function and refining the distillation learning protocol (Gu et al., 2023; Berthelot et al., 2023; Heek et al., 2024), resulting in improved results.
+Parameter-Efficient Fine-Tuning.
+Training a diffusion model is highly resource-intensive and environmentally unfriendly.
+Fine-tuning such a model can also be challenging due to the vast number of parameters involved (Aghajanyan et al., 2020).
+Thus, Parameter-Efficient Fine-Tuning (PEFT) (Houlsby et al., 2019) was proposed to enable the fine-tuning of pretrained models with a limited number of parameters required for training.
+Among these techniques, Low-Rank Adaptation (LoRA) (Hu et al., 2021) has demonstrated transcendental performance.
+LoRA’s strategy involves freezing the pretrained model weights and injecting trainable rank decomposition matrices, which succinctly represent the required adjustments in the model’s weights for fine-tuning.
+With this strategy, LoRA significantly reduces the volume of parameters to be modified, thereby substantially decreasing both computational load and storage demands.
 3 Preliminaries
@@
 Denote the marginal distribution of $\bm{x}_{t}$ following the forward process as $p_{t}(\bm{x}_{t})$ and, such an Itô SDE gradually perturbs the empirical data distribution $p_{0}(\bm{x})=p_{\rm{data}}(\bm{x})$ towards the prior distribution $p_{T}(\bm{x})\approx\pi(\bm{x})$ approximately, where $\pi(\bm{x})$ is a tractable Gaussian distribution.
-(a) Training process, wherein the TCF expands the boundary conditions to an arbitrary timestep of $s$, thereby reducing the theoretical upper limit of error. The details can be found in Algorithm 1.
-(b) Sampling process, as detailed in Algorithm 3 & Algorithm 4, where SSS significantly reduces accumulated error through the iterative traversal with the stochastic parameter $\gamma$ compared to the multistep consistency sampling.
-Figure 2: The comparative overview of the baseline Consistency Distillation (Song et al., 2023) and the proposed Trajectory Consistency Distillation, includes Trajectory Consistency Function (TCF) for training and Strategic Stochastic Sampling (SSS) for inference.
 Remarkably, Song et al.
@@
 There already exist off-the-shelf ODE solvers (Song et al., 2020b, a; Karras et al., 2022) or efficient numerical solvers (Lu et al., 2022a, b; Zhang & Chen, 2022) that can be directly applied to approximate the exact solution.
+(a) Training process, wherein the TCF expands the boundary conditions to an arbitrary timestep of $s$, thereby reducing the theoretical upper limit of error. The details can be found in Algorithm 1.
+(b) Sampling process, as detailed in Algorithm 3 & Algorithm 4.
+Figure 2: The comparative overview of the baseline Consistency Distillation (Song et al., 2023) and the proposed Trajectory Consistency Distillation, includes Trajectory Consistency Function (TCF) for training and Strategic Stochastic Sampling (SSS) for inference.
 3.2 Consistency Models
@@
 Typically, $\bm{f}$ can be distilled from a pretrained diffusion model $F_{\bm{\theta}}(\bm{x}_{t},t)$ and parameterized as:
-$\bm{f}_{\bm{\theta}}(\tilde{\bm{x}}_{t},t)=\begin{cases}\bm{x}_{0},&t=0\\ \texttt{Solver}(F_{\bm{\theta}}(\tilde{\bm{x}}_{t},t),t,0;\bm{\theta}),&t\in(0,T]\end{cases}$
-The $\texttt{Solver}(\cdot,t,0;\theta)$ is the update function of an ODE solver used to estimate ${\bm{x}}_{0}$ when given the output $F_{\theta}(\tilde{\bm{x}}_{t},t)$ from a pretrained model at timestep $t$.
+$\bm{f}_{\bm{\theta}}(\bm{x}_{t},t)=\begin{cases}\bm{x}_{0},&t=0\\ F_{\bm{\theta}}(\bm{x}_{t},t),&t\in(0,T]\end{cases}$ (6)
+or using skip-connection as:
+${\bm{f}}_{\bm{\theta}}({\bm{z}}_{t},{\bm{c}},t)=c_{\text{skip}}(t){\bm{z}}_{t}+c_{\text{out}}(t)F_{\bm{\theta}}(\bm{x}_{t},t),$ (7)
+where $c_{\text{skip}}(\epsilon)=1$ and $c_{\text{out}}(\epsilon)=0$.
 For training consistency models, the objective of Consistency Distillation (CD) is defined as minimizing:
@@
 where $0=t_{1}<t_{2}\cdots<t_{N}=T$, $n$ uniformly distributed over $\{1,2,\cdots,N-1\}$, $\lambda(\cdot)\in\mathbb{R}^{+}$ is a positive weighting function, and the expectation is taken with respect to $\bm{x}\sim p_{\rm{data}}$.
-$\bm{x}_{t_{n+1}}$ can be sampled using SDE  1 and $\hat{\bm{x}}_{t_{n}}^{\bm{\phi}}$ is calculated by $\hat{\bm{x}}_{t_{n}}^{\bm{\phi}}:=\Phi(\bm{x}_{t_{n+1}},t_{n+1},t_{n};\bm{\phi})$, where $\Phi(\cdots;\bm{\phi})$ represents the update function of a one-step ODE solver applied to the empirical PF ODE 3.
+$\bm{x}_{t_{n+1}}$ can be sampled using SDE (1) and $\hat{\bm{x}}_{t_{n}}^{\bm{\phi}}$ is calculated by $\hat{\bm{x}}_{t_{n}}^{\bm{\phi}}:=\Phi(\bm{x}_{t_{n+1}},t_{n+1},t_{n};\bm{\phi})$, where $\Phi(\cdots;\bm{\phi})$ represents the update function of a one-step ODE solver applied to the empirical PF ODE (3).
 Moreover, $\bm{\theta^{-}}$ is introduced to stabilize the training process and updated by an exponential moving average (EMA) strategy, i.e., given $10\leq\mu<$ 1, $\bm{\theta}^{-}\leftarrow\texttt{sg}(\mu\bm{\theta}^{-}+(1-\mu)\bm{\theta})$.
+With a well-trained model $\bm{\theta^{*}}$ such that $\mathcal{L}_{\rm{CD}}^{N}(\bm{\theta^{*}},\bm{\theta^{*}};\bm{\phi})=0$, and considering that the ODE solver $\Phi(\cdots;\bm{\bm{\phi}})$ has the local discretisation error, Theorem 1 in (Song et al., 2023) shows that the consistency distillation error is bounded:
+$\mathop{\textnormal{sup}}\limits_{n,{\bm{x}}}\|{\bm{f}}_{\bm{\theta^{*}}}({\bm{x}},t_{n})-{\bm{f}}({\bm{x}},t_{n};\bm{\phi})\|_{2}=\mathcal{O}\left((\Delta t)^{p}\right),$ (9)
+with $\Delta t$ and $p$ as defined in (Song et al., 2023).
 Besides the distillation strategy that needs an existing score model, Song et al.
@@
 In this paper, our primary focus is on the CD objective.
-3 Elucidating Errors in Consistency Models
-In this section, we elucidate various types of errors that occurred in multistep sampling of consistency models, which motivates us to propose corresponding solutions aimed at refining them.
-Our analysis reveals that the error is primarily composed of three components: distillation error of consistency models, estimation error of score matching models, and accumulated discretisation error (also known as truncation error) during solving the ODEs.
-3.1 Consistency Distillation Error
-With a well-trained model $\bm{\theta^{*}}$ such that $\mathcal{L}_{\rm{CD}}^{N}(\bm{\theta^{*}},\bm{\theta^{*}};\bm{\phi})=0$, and considering that the ODE solver $\Phi(\cdots;\bm{\bm{\phi}})$ has the local discretisation error, Theorem 1 in (Song et al., 2023) shows that the consistency distillation error is bounded:
-$\mathop{\textnormal{sup}}\limits_{n,{\bm{x}}}\|{\bm{f}}_{\bm{\theta^{*}}}({\bm{x}},t_{n}),{\bm{f}}({\bm{x}},t_{n};\bm{\phi})\|_{2}=\mathcal{O}\left((\Delta t)^{p}\right),$ (7)
-with $\Delta t$ and $p$ as defined in (Song et al., 2023).
-3.2 Error Bound in Multistep Consistency Sampling
-Theoretically, with a well-trained consistency model ${\bm{f}}_{\bm{\theta^{*}}}$, one can generate samples with just one forward pass through the consistency model.
-However, the one-step sampling yields suboptimal results (Luo et al., 2023a; Song et al., 2023).
-Thus, multistep consistency sampling has been introduced in (Song et al., 2023) to enhance sample quality through alternating denoising and noise injection steps.
-We term it as multistep sampling in this manuscript.
-For simplicity, we consider VP SDEs in this paper; thus, the drift and diffusion coefficients $\mu(t)$ and $\nu(t)$ in SDE( Equation 1) can be written as:
-$\mu(t)=\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t},\quad\nu(t)=\sqrt{\frac{\mathrm{d}\sigma_{t}^{2}}{\mathrm{d}t}-2\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}\sigma_{t}^{2}},$ (8)
-where $\alpha_{t}$ and $\sigma_{t}$ specify the noise schedule in the perturbation kernels,
-$q_{0t}({\bm{x}}_{t}|{\bm{x}}_{0})=\mathcal{N}({\bm{x}}_{t}|\alpha_{t}{\bm{x}}_{0},\sigma_{t}^{2}\bm{I}).$ (9)
-Given a sequence of $N$ sampling timesteps $T=\tau_{1}>\tau_{2}>\cdots>\tau_{N}$ and an initial value $\tilde{{\bm{x}}}_{\tau_{1}}\sim\mathcal{N}(0,\bm{I})$, the generating procedure of the $n$-th step can be written as:
-${\bm{x}}_{\tau_{1}\rightarrow 0}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}(\tilde{{\bm{x}}}_{\tau_{1}},T),$ (10)
-Diffuse: $\tilde{{\bm{x}}}_{\tau_{n}}\leftarrow\alpha_{\tau_{n}}{\bm{x}}_{\tau_{(n-1)}\rightarrow 0}+\sigma_{\tau_{n}}{\bm{z}},\quad{\bm{z}}\sim\mathcal{N}(0,\bm{I})$
-Denoise: ${\bm{x}}_{\tau_{n}\rightarrow 0}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}(\tilde{{\bm{x}}}_{\tau_{n}},\tau_{n}),$
-this process is detailed in Algorithm 3. We derive a further corollary from Corollary 7 in  (Lyu et al., 2023) as follows:
-Corollary 3.1.
-With the sampling process defined in Equation 10 and denoting the distribution of ${\bm{x}}_{\tau_{n}\rightarrow 0}$ as $q_{\bm{\theta^{*}},n}$, we have single-step sampling result, $q_{\bm{\theta^{*}},1}={\bm{f}}_{{\bm{\theta^{*}}},T}\sharp\mathcal{N}(0,\bm{I})$ and multistep sampling result, $q_{\bm{\theta^{*}},N}={\bm{f}}_{{\bm{\theta^{*}}},\tau_{N}}\sharp\mathcal{N}(\alpha_{\tau_{N}}{\bm{x}}_{\tau_{(N-1)}\rightarrow 0},\sigma_{\tau_{N}}\bm{I})$ with $N>1$.
-The Total Variational (TV) distance between $q$ and $p_{\rm{data}}$ is
-$TV(q_{\bm{\theta^{*}},1},p_{\rm{data}})$ $=\mathcal{O}\left(T(\varepsilon_{cd}+\mathcal{L}_{f}\varepsilon_{se})\right)$
-$TV(q_{\bm{\theta^{*}},N},p_{\rm{data}})$ $=\mathcal{O}\left({2^{-N}T(\varepsilon_{cd}+\mathcal{L}_{f}\varepsilon_{se})}\right),$
-where normal-$\sharp$ is a push-forward operator associated with a measurable map, $\mathcal{L}_{f}$ is the Lipschitz constant of consistency model, and $\varepsilon_{cd}$, $\varepsilon_{se}$ represent the error of consistency distillation in Equation 7 and score matching estimation, respectively.
-The detailed proof can be found in Appendix D.1.
-From  Corollary 3.1, we observe that multistep sampling has a lower error bound than a single step.
-This observation aligns with the empirical findings presented in (Karras et al., 2022): the local truncation error scales superlinearly with respect to step size, thus increasing $N$ improves the accuracy of the solution.
-Moreover, it illustrates that the model performance tends to underperform single-step generation methods (Sauer et al., 2023; Yin et al., 2023) without the help of additional supervision.
-3.3 Accumulated Error in Multistep Sampling
-However, due to the error of estimation and discretisation in practice, the total error will be accumulated in every sampling step.
-Denote the distribution of ${\bm{x}}_{\tau_{n}\rightarrow 0}$ output by the true consistency function ${\bm{f}}(\cdot,\cdot;\bm{\phi})$ as $p_{n}$, Theorem 2 in (Chen et al., 2022) demonstrates that the TV error between $q_{\bm{\theta^{*}},n}$ and $p_{n}$ is
-$TV(q_{\bm{\theta^{*}},n},p_{n})=\mathcal{O}(\sqrt{\tau_{n}}),$ (11)
-the error in each $n$-th step will accumulate over $N$ sampling steps, resulting in the defined accumulated error
-$TV(q_{\bm{\theta^{*}},N},p_{N})=\mathcal{O}\left(\sum_{n=1}^{N}\sqrt{\tau_{n}}\right).$ (12)
-The above result is a special case of Theorem 4.2, and its proof is presented in Section D.3. As a consequence, there is a significant reduction in image details at higher NFEs, as shown in Figure 3.
+3.3 Latent Consistency Model
+The Latent Consistency Model successfully extends consistency models (Song et al., 2023) to the image latent space corresponding with an auto-encoder (Rombach et al., 2022).
+To leverage the advantages of CMs, the authors proposed Latent Consistency Distillation, where the consistency function with condition information ${\bm{c}}$ can be defined as:
+$\bm{f}(\bm{z}_{t},{\bm{c}},t)\mapsto\bm{z}_{0},$ (10)
+which can directly predict the solution of PF-ODE for $t=0$.
+To leverage the knowledge of pretrained text-to-image models effectively, LCMs introduce the parameterisation method when parameterizing the consistency model with $\epsilon$-prediction teacher model:
+$\tilde{{\bm{f}}}_{\bm{\theta}}({\bm{z}}_{t},{\bm{c}},t)=c_{\text{skip}}(t){\bm{z}}_{t}+c_{\text{out}}(t)\left(\frac{{\bm{z}}_{t}-\sigma_{t}\hat{\epsilon}_{\bm{\theta}}({\bm{z}}_{t},{\bm{c}},t)}{\alpha_{t}}\right),$ (11)
+where the boundary condition satisfies $c_{\text{skip}}(0)=1$ and $c_{\text{out}}(0)=0$, $\alpha_{t}$ and $\sigma_{t}$ specify the noise schedule in the perturbation kernels.
+It’s worth noting that the above function, with appropriate parameterisation, is compatible with different prediction methods (e.g., $x$-prediction, $\epsilon$-prediction in (Ho et al., 2020), $v$-prediction in (Salimans & Ho, 2022)).
+To accelerate distillation training, Luo et al.
+(2023a) proposed the skipping-step method to ensure self-consistency between timesteps $k$-steps apart, $t_{n+k}\rightarrow t_{n}$, instead of the original adjacent timesteps, $t_{n+1}\rightarrow t_{n}$.
+Additionally, Luo et al.
+(2023a) further implemented one-stage guided distillation (Meng et al., 2023) by solving the augmented PF-ODE.
+Then, the solution of $t_{n}$ from $t_{n+k}$ can be approximated with DDIM (Song et al., 2020a) or DPM-Solver (Lu et al., 2022a, b):
+$\begin{split}\hat{{\bm{z}}}_{t_{n}}^{\phi,\omega}=(1+\omega)&\Phi({\bm{z}}_{t_{n+k}},t_{n+k},t_{n},{\bm{c}};\bm{\phi})\\ &-\omega\Phi({\bm{z}}_{t_{n+k}},t_{n+k},t_{n},\varnothing;\bm{\phi}),\end{split}$ (12)
+where $\Phi$ represents the update function of the PF-ODE solver with the pre-trained parameter $\phi$, and $\omega$ is the classifier-free guidance weight sampled from $\omega\in[\omega_{\text{min}},\omega_{\text{max}}]$.
+The objective function of LCM is derived from the self-consistency property, similar to the consistency distillation loss in (Song et al., 2023):
+$\mathcal{L}_{\rm{LCM}}(\bm{\theta},$ $\bm{\theta^{-}};\bm{\Phi}):=\mathbb{E}_{{\bm{z}},{\bm{c}},\omega,n}\bigg{[}$ (13)
+$d\left(\tilde{\bm{f}}_{\bm{\theta}}({\bm{z}}_{t_{n+k}},{\bm{c}},t_{n+k});\tilde{\bm{f}}_{\bm{\theta}^{-}}(\hat{{\bm{z}}}_{t_{n}}^{\Phi,\omega},{\bm{c}},t_{n})\right)\bigg{]},$
+where $d$ represents the distance between samples, with a common choice being the $\ell_{2}$ distance.
 4 Trajectory Consistency Distillation
-4.1 Trajectory Consistency Function
-Definition.
-The distillation error from the consistency model can be reduced by expanding the original boundary conditions to encompass the entire trajectory.
-To facilitate this, we introduce the Trajectory Consistency Function (TCF), designed to enable comprehensive tracking along the full trajectory.
-${\bm{f}}({\bm{x}}_{t},t,s)\mapsto{\bm{x}}_{s}.$ (13)
-The trajectory consistency function possesses the property of trajectory consistency, enhancing the original self-consistency property in an endpoint-unrestricted trajectory.
-Specifically, its output remains consistent for arbitrary sets $({\bm{x}}_{t},t)$ with the given $s$ that belong to the same PF ODE trajectory, where $0\leqslant s\leqslant t\leqslant T$,
-${\bm{f}}({\bm{x}}_{t},t,s)={\bm{f}}({\bm{x}}^{\prime}_{t},t^{\prime},s)\quad\forall t,t^{\prime},s\in[0,T].$ (14)
+4.1 Background
+Given the $\epsilon$-prediction model $\epsilon_{\bm{\theta}}(\cdot)$, Luo et al.
+(2023a) used DDIM (Song et al., 2020a) to re-parameterize the consistency function in (Song et al., 2023) but still followed the skip connection form as in Equation 11 with boundary conditions $c_{\text{skip}}(0)=1$ and $c_{\text{out}}(0)=0$.
+However, such a parameterization process includes a discretization error caused by the finite difference approximation of an explicit numerical method (Song et al., 2020a) with the spacing $t\rightarrow 0$.
+Besides, in practical experiments with LCM released training codes  [^1] https://github.com/luosiallen/latent-consistency-model, we noticed the implementation of preconditions $c_{\text{skip}}$ and $c_{\text{out}}$ only matches the boundary condition but stray from the initial definition in (Karras et al., 2022).
+In order to better fit the formulation in LDM (Rombach et al., 2022) and SDXL (Podell et al., 2023), we first dispose of the skip-connection form with preconditions $c_{\text{skip}}$ $c_{\text{out}}$ and rewrite Equation 11 as:
+$\begin{split}\tilde{{\bm{f}}}_{\bm{\theta}}({\bm{x}}_{t},t)&=\frac{{\bm{x}}_{t}-\sigma_{t}\hat{\epsilon}_{\bm{\theta}}({\bm{x}}_{t},t)}{\alpha_{t}}\\ &=\frac{1}{\alpha_{t}}{\bm{x}}_{t}-\frac{\sigma_{t}}{\alpha_{t}}\hat{\epsilon}_{\bm{\theta}}(\hat{{\bm{x}}}_{t},t)\\ &=\frac{\alpha_{0}}{\alpha_{t}}{\bm{x}}_{t}-\alpha_{0}\left(\frac{\sigma_{t}}{\alpha_{t}}-\frac{\sigma_{0}}{\alpha_{0}}\right)\hat{\epsilon}_{\bm{\theta}}(\hat{{\bm{x}}}_{t},t),\end{split}$ (14)
+which is in the form of DDIM (Song et al., 2020a). As we mentioned before, this function also includes an approximation error with huge spacing.
+Lu et al.
+(2022a) pointed out Equation 14 is actually an approximated case for the exact solution of PF-ODEs with semi-linear structure with an exponential integrator in terms of $\bm{\epsilon}_{\bm{\theta}}$:
+$\hat{\bm{x}}_{s}=\frac{\alpha_{s}}{\alpha_{t}}{\bm{x}}_{t}+\alpha_{s}\int_{\lambda_{t}}^{\lambda_{s}}e^{-\lambda}\hat{\bm{\epsilon}}_{\bm{\theta}}(\hat{{\bm{x}}}_{\lambda},\lambda)\text{d}\lambda,$ (15)
+and analyzed the corresponding error associated with $\lambda_{t}$ and $\lambda_{s}$, characterized as $\mathcal{O}((\lambda_{s}-\lambda_{t})^{k})$ with $k$-th Taylor expansion.
+In this context, $\bm{\epsilon}_{\bm{\theta}}$ is a trainable network with parameter $\bm{\theta}$, and $\lambda_{t}:=\log(\alpha_{t}/\sigma_{t})$ is the log-SNR.
+We note that when using the same order $k$ of Taylor expansion, the error decreases as the interval between $s$ and $t$ reduces.
+Another point worth noting is that shortening the time intervals makes the boundedness assumption in (Lu et al., 2022a) easier to meet, which ensures the error correspondingly decreases as the order $k$ increases.
+We would like to highlight that such an mentioned error is the discretization error arising during the parameterisation process in  Equation 11.
+To further improve the guided sampling quality, another exact solution of PF-ODEs with ${\bm{x}}_{\bm{\theta}}$ is proposed in (Lu et al., 2022b):
+$\hat{{\bm{x}}}_{s}=\frac{\sigma_{s}}{\sigma_{t}}{\bm{x}}_{t}+\sigma_{s}\int_{\lambda_{t}}^{\lambda_{s}}e^{\lambda}\hat{{\bm{x}}}_{\bm{\theta}}(\hat{{\bm{x}}}_{\lambda},\lambda)\text{d}\lambda,$ (16)
+which leads to better guided sampling performance and shares analogous properties with Equation 15.
+However, given that the consistency function is devoid of any guidance sampling, these two equations are equally valid views and can be used interchangeably (Kingma et al., 2021).
+In this work, we opt for the ${\bm{x}}_{\bm{\theta}}$ representation due to its greater intuitiveness, despite the fact that $\bm{\epsilon}_{\bm{\theta}}$ is mostly used for parameterization in practice.
+Furthermore, Equation 9 also reveals that in the distillation process, the upper bound of the CD loss is not only related to the number of steps $\Delta t$ derived by solvers, but is also associated with the time interval $t\rightarrow 0$ spanned in the consistency function.
 Figure 3: Synthesis results across various NFEs.
 Due to accumulated errors in multistep sampling, LCM experiences a loss of image detail, leading to a degradation in performance, whereas TCD addresses this issue.
-Additional samples are available in Section E.1.
+Additional samples are available in Section D.1.
+4.2 Trajectory Consistency Function
+Such errors as mentioned lead us to seek a better form for the consistency function by broadening the endpoint from the origin to a midpoint $s$ between 00 and $t$ within the semi-linear structure.
+Recall Equation 6 and the definition of the consistency function in (Song et al., 2023): ${\bm{f}}({\bm{x}}_{t},t)\mapsto{\bm{x}}$, and the definition of the self-consistency property of the consistency function  Equation 5: The outputs of the consistency function are consistent for arbitrary pairs that belong to the same PF-ODE trajectory.
+The goal of consistency distillation is to estimate this consistency function and ensure the self-consistency property.
+Definition.
+We introduce the Trajectory Consistency Function (TCF) to re-parameter the output of ${\bm{x}}_{\bm{\theta}}(\cdot)$ to ${\bm{x}}_{s}$:
+${\bm{f}}^{\rightarrow s}({\bm{x}}_{t},t)\mapsto{\bm{x}}_{s}.$ (17)
+Regarding the selection of $s$, our objective is to ensure that $s$ conforms to the minimum error across various discrete situations; thus, $s$ is designated as an arbitrary integer within the range $[0,t]$.
+And TCF also supports self-consistency. TCF remains consistent for arbitrary sets $({\bm{x}}_{t},t)$ with the given $s$ that belong to the same PF ODE trajectory:
+${\bm{f}}^{\rightarrow s}({\bm{x}}_{t},t)={\bm{f}}^{\rightarrow s}({\bm{x}}_{t^{\prime}},t^{\prime})\quad\forall t,t^{\prime},s\in[0,T].$ (18)
 Parameterisation.
-The semi-linear structure of the empirical PF-ODE revealed by (Lu et al., 2022a, b; Zhang & Chen, 2022) motivates us to parameterise the trajectory consistency function using the exponential integrators form, as shown in Equation 15.
-${\bm{f}}_{\bm{\theta}}({\bm{x}}_{t},t,s)=\frac{\sigma_{s}}{\sigma_{t}}{\bm{x}}_{t}+\sigma_{s}\int_{\lambda_{t}}^{\lambda_{s}}e^{\lambda}\hat{{\bm{x}}}_{\bm{\theta}}(\hat{{\bm{x}}}_{\lambda},\lambda)\text{d}\lambda,$ (15)
-where $\lambda_{t}:=\log(\alpha_{t}/\sigma_{t})$ is the log-SNR, and ${\bm{x}}_{\bm{\theta}}$ is a trainable network with parameter $\bm{\theta}$.
+The semi-linear structure with the exponential integrators form (Lu et al., 2022a, b; Zhang & Chen, 2022) can be easily used to parameterise the trajectory consistency function,
+[…]
 For $k\geqslant 1$, one can take the ($k$-1)-th Taylor expansion at $\lambda_{t}$ for ${\bm{x}}_{\bm{\theta}}$ w.r.t $\lambda\in[\lambda_{s},\lambda_{t}]$, we have:
-$\begin{split}{\bm{f}}_{\bm{\theta}}&({\bm{x}}_{t},t,s)=\frac{\sigma_{s}}{\sigma_{t}}{\bm{x}}_{t}+\\ &\sigma_{s}\sum_{n=0}^{k-1}{\bm{x}}_{\bm{\theta}}^{(n)}(\hat{{\bm{x}}}_{\lambda_{t}},\lambda_{t})\int_{\lambda_{t}}^{\lambda_{s}}e^{\lambda}\frac{(\lambda-\lambda_{t})^{n}}{n!}\text{d}\lambda+\mathcal{O}(h^{k+1}),\end{split}$ (16)
+$\begin{split}{\bm{f}}^{\rightarrow s}_{\bm{\theta}}&({\bm{x}}_{t},t)=\frac{\sigma_{s}}{\sigma_{t}}{\bm{x}}_{t}+\\ &\sigma_{s}\sum_{n=0}^{k-1}\hat{\bm{x}}_{\bm{\theta}}^{(n)}(\hat{{\bm{x}}}_{\lambda_{t}},\lambda_{t})\int_{\lambda_{t}}^{\lambda_{s}}e^{\lambda}\frac{(\lambda-\lambda_{t})^{n}}{n!}\text{d}\lambda+\mathcal{O}(h^{k+1}),\end{split}$ (20)
 where $h=\lambda_{s}-\lambda_{t}$ and ${\bm{x}}_{\bm{\theta}}^{(n)}(\cdot,\cdot)$ is the $n$-th order total derivatives of ${\bm{x}}_{\bm{\theta}}$ w.r.t $\lambda$.
-Here, we consider 1st-order and 2nd-order estimations, omitting high-order terms $\mathcal{O}(h^{k+1})$.
+Here, we consider 1st-order and 2nd-order estimations, omitting high-order terms $\mathcal{O}(h^{k+1})$ in equation expression.
 […]
 For the $2$nd-order expansion, we can write the trajectory consistency function as:
-$\begin{split}\text{TCF(2):}\quad{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t},&t,s)=\frac{\sigma_{s}}{\sigma_{t}}{\bm{x}}_{t}-\alpha_{s}(e^{-h}-1)\\ &\left((1-\frac{1}{2r})\hat{\bm{x}}_{\theta}({\bm{x}}_{t},t)+\frac{1}{2r}\hat{\bm{x}}_{\theta}(\hat{\bm{x}}_{u},u)\right),\end{split}$ (18)
+$\begin{split}\text{TCF(2):}\quad\tilde{{\bm{f}}}^{\rightarrow s}_{\bm{\theta}}({\bm{x}}_{t},&t)=\frac{\sigma_{s}}{\sigma_{t}}{\bm{x}}_{t}-\alpha_{s}(e^{-h}-1)\\ &\left((1-\frac{1}{2r})\hat{\bm{x}}_{\theta}({\bm{x}}_{t},t)+\frac{1}{2r}\hat{\bm{x}}_{\theta}(\hat{\bm{x}}_{u},u)\right),\end{split}$ (22)
 where $u$ is the mid-timestep w.r.t. $t>u>s$ and $r:=(\lambda_{u}-\lambda_{t})/h$.
+According to Theorem 3.2 in (Lu et al., 2022a), the upper error bounds for TCF parameterisation is the same as DPM-Solver-$k$.
+Let ${\bm{f}}_{\bm{\theta}}^{\rightarrow s}({\bm{x}}_{t},t)$ be the ideal trajectory consistency function with exponential integrators form of the empirical PF-ODE, where $t$ and $s$ are the start timestep and end timestep along PF-ODE, we have the discretisation error regarding to the finite $k$-th order of TCF:
+$\|\tilde{{\bm{f}}}_{\bm{\theta}}^{\rightarrow s}({\bm{x}}_{t},t)-{\bm{f}}_{\bm{\theta}}^{\rightarrow s}({\bm{x}}_{t},t)\|_{2}=\mathcal{O}(h_{\Delta t}^{k+1}),$ (23)
+where $h:=\lambda_{s}-\lambda_{t}$ and $k$ is the TCF order. As identified in  (Zhang & Chen, 2022; Lu et al., 2022a), employing semi-linearity in diffusion ODEs results in minimal error.
 We also propose a modified network $F_{\bm{\theta}}$ with additional parameters for conditioning on $s$ to directly estimate the exponentially weighted integral of ${\bm{x}}_{\bm{\theta}}$ without omitting residual term when $k=1$ in Equation 20:
 […]
-Broadening the Boundary Condition.
+In order to unify the representation of TCF, we abuse
+${\bm{f}}({\bm{x}}_{t},t,s):={\bm{f}}^{\rightarrow s}({\bm{x}}_{t},t),$ (25)
+where $s$ can be a parameter either in the reparameterised process or the network.
+We observe that TCF(S+) integrates the parameter $s$ into the network, yielding a formulation akin to the $G$ function in  (Kim et al., 2023), albeit with a different coefficient.
+The underlying reason is that both instances employ a neural network to directly get the integral term from $t$ to $s$.
+The key distinction lies in the dependency of CTM’s $G_{\theta}$ on the Euler Solver and the requirement of a ${\bm{g}}_{\theta}$ for estimating $\mathbb{E}[{\bm{x}}|{\bm{x}}_{t}]$.
+Conversely, TCFs adopt a semi-linear structure for a better consistency function and use the finite difference approximation to parameterize the integral term.
+Broaden Boundary Conditions
 The boundary condition in the CM is limited to intervals extending from any start point on the solution trajectory to the origin.
-Conversely, TCF mitigates this constraint, allowing the model to handle any interval along the PF-ODE trajectory and preventing trivial solutions ${\bm{f}}_{\bm{\theta}}({\bm{x}}_{t},t,s)\equiv 0$ from arising in TCD training.
-Consequently, we broaden the boundary condition to encompass a more comprehensive range of trajectory intervals,
-${\bm{f}}_{\bm{\theta}}({\bm{x}}_{s},s,s)={\bm{x}}_{s},$ (20)
-it is evident that all three types of our parametrisation satisfy the broad boundary condition effortlessly.
+And TCF is required to adhere to classical boundary conditions (Song et al., 2020b)
+${\bm{f}}({\bm{x}}_{0},0,0)={\bm{x}}_{0},$ (26)
+but also needs to meet a broaden boundary conditions
+${\bm{f}}({\bm{x}}_{s},s,s)={\bm{x}}_{s},$ (27)
+to encompass a more comprehensive range of trajectory intervals with arbitrary sets.
+It is evident that our parametrisation satisfy both the classical boundary conditions and the broaden boundary condition effortlessly.
 Training.
@@
 $\hat{{\bm{x}}}_{t_{n}}^{\phi,k}=\Phi^{(k)}({\bm{x}}_{t_{n+k}},t_{n+k},t_{n};\bm{\phi}).$ (28)
-Thus, we could express the object of trajectory distillation in alignment with reconstruction:
+Thus, we could express the object of trajectory distillation in alignment with reconstruction, similar to the consistency distillation loss in (Song et al., 2023):
 $\mathcal{L}^{N}_{\rm{TCD}}(\bm{\theta},$ $\bm{\theta^{-}};\bm{\phi}):=\mathbb{E}[\omega(t_{n},t_{m})$ (29)
@@
 The detailed training process is outlined in Algorithm 1.
-Below, we offer a theoretical justification based on asymptotic analysis to explain how trajectory consistency distillation optimizes the distillation error.
-Theorem 4.1.
-Let $\Delta t:=\rm{max}_{n\in[\![1,N-1]\!]}\{\lvert t_{n+1}-t_{n}\rvert\}$, and $\bm{f}(\cdot,\cdot,\cdot;\bm{\phi})$ be the trajectory consistency function of the empirical PF ODE in Equation 3.
-Assume $\bm{f}_{\bm{\theta}}$ satisfies the Lipschitz condition, that is, there exists $L>0$ such that for all $t\in[0,T]$, $\bm{x}$ and $\bm{y}$, we have $\left\lVert{\bm{f}_{\bm{\theta}}(\bm{x},t,s)-\bm{f}_{\bm{\theta}}(\bm{y},t,s)}\right\rVert_{2}\leqslant L\left\lVert{\bm{x}-\bm{y}}\right\rVert_{2}$.
-Assume further that for all $n\in[\![1,N-1]\!]$, the $p$-th order ODE solver called at $t_{n+1}$ has local error uniformly bounded by $\mathcal{O}((t_{n+1}-t_{n})^{p+1})$ with $p\geqslant 1$.
-Then, if there is a $\bm{\theta^{*}}$ so that $\mathcal{L}_{\rm{TCD}}^{N}(\bm{\theta^{*}},\bm{\theta^{*}};\bm{\phi})=0$, for any $n\in[\![1,N-1]\!]$ and $m\in[\![1,n]\!]$, we have
-$\begin{split}\mathop{\textnormal{sup}}\limits_{n,m,{\bm{x}}}\|{\bm{f}}_{\bm{\theta^{*}}}({\bm{x}},t_{n},t_{m}),{\bm{f}}({\bm{x}},&t_{n},t_{m};\phi)\|_{2}\\ &=\mathcal{O}\left((\Delta t)^{p}\right)(t_{n}-t_{m}).\end{split}$
-Proof.
-The proof is provided in  Section D.2. ∎
-Theorem 4.1 implies that the distillation error of TCF is upper bound by that of CD presented in Equation 7.
+This objective function inherit the form in (Luo et al., 2023a; Song et al., 2023), we expand it to a broaden condition according to the definition of TCF in Equation 17 and ${\bm{f}}_{\bm{\theta}}({\bm{x}}_{t},t,s)$ still holds the form of ${\bm{f}}^{\rightarrow s}_{\bm{\theta}}({\bm{x}}_{t},t)$, which can be easily reparameterised and lead to a fast convergence.
+Such parameterisation of consistency function in semi-linear is pivotal for fast convergence and better results.
+From Theorem 1 in (Song et al., 2023), we could found trajectory consistency distillation with a shorter spacing can reduce the distillation error from $\mathcal{O}\left((\Delta t)^{p}\right)(t_{n}-t_{0})$ to $\mathcal{O}\left((\Delta t)^{p}\right)(t_{n}-t_{m})$.
+It is observed that the TCF not only leads the reduction in the parameterisation error bound as delineated in Equation 23, but also lowers the distillation error bound.
 4.3 Strategic Stochastic Sampling
-The proposed trajectory consistency function not only optimizes the distillation loss but also enables the model to access non-origin destinations along the PF ODE.
-This capability allows Strategic Stochastic Sampling (SSS) to further reduce discretisation errors and estimation errors introduced in each sampling step.
+It is worth noting that once our TCF is trained, one can use common PF-ODE solvers like DDIM or SDE-DPM-Solver to sample high quality images.
+But it is not easy to control the stochastic strength with existing solvers; hence, we proposed Strategic Stochastic Sampling for better stochasticity control.
+Background.
+Recalling the Multistep Consistency Sampling process in Algorithm 3, each iteration of Multistep Consistency Sampling from $t$ to $s$ can be expressed as:
+Denoise: ${\bm{x}}_{t\rightarrow 0}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}({{\bm{x}}}_{t},t),$ (30)
+Diffuse: $\hat{{\bm{x}}}_{s}\leftarrow\alpha_{s}{\bm{x}}_{t\rightarrow 0}+\sigma_{s}{\bm{z}},$
+we found that the above process with parameterisation consistency function (Equation 14) can be rewritten as a DDIM sampling equation  (Song et al., 2020a) with $\eta=\sqrt{1-\alpha_{s}^{2}}$ in Equation 32:
+${\bm{x}}_{s}=\alpha_{s}\underbrace{\left(\frac{{\bm{x}}_{t}-\sigma_{t}\epsilon_{\bm{\theta}}({\bm{x}}_{t})}{\alpha_{t}}\right)}_{\text{denoise}}+\underbrace{\sigma_{s}{\bm{z}}}_{\text{diffuse}}.$ (31)
+Equation 31 leads us to review the general version of DDIM:
+$\begin{split}{\bm{x}}_{s}&=\alpha_{s}{\left(\frac{{\bm{x}}_{t}-\sigma_{t}\epsilon_{\bm{\theta}}({\bm{x}}_{t})}{\alpha_{t}}\right)}\\ &\quad\quad\quad\quad+\sqrt{1-\alpha_{s}^{2}-\eta^{2}}\epsilon_{\bm{\theta}}({\bm{x}}_{t})+\eta{\bm{z}},\end{split}$ (32)
+where occurs a parameter $\eta$ to control the stochasticity.
+And many related works (Karras et al., 2022; Song et al., 2020a; Zhang & Chen, 2022) observed that few stochasticity can help reduce the sampling error, but high stochasticity will lead to poor quality.
+Such stochasticity during multistep sampling will obviously degrade the final output.
+Inspired by the ideas of stochasticity control term from DDIM and SDE-DPM-Solvers (Lu et al., 2022b), we propose Strategic Stochastic Sampling (SSS), which explicitly integrates the noise control term as an input parameter in the sampling process.
+Besides, our SSS could also be extended to different orders based on SDE-DPM-Solvers (Lu et al., 2022b).
+Here we demonstrate the 1-st order SSS solver, which is based on Equation 32:
+$\begin{split}&{\bm{x}}_{s}=\alpha_{s}\left(\frac{{\bm{x}}_{t}-\sigma_{t}\epsilon_{\bm{\theta}}({\bm{x}}_{t})}{\alpha_{t}}\right)+\alpha_{s}\frac{\sigma_{s^{{}^{\prime}}}}{\alpha_{s^{{}^{\prime}}}}\epsilon_{\bm{\theta}}({\bm{x}}_{t})+\sqrt{1-\frac{\alpha_{s}^{2}}{\alpha_{s^{{}^{\prime}}}^{2}}}{\bm{z}}\\ &=\frac{\alpha_{s}}{\alpha_{s^{{}^{\prime}}}}\underbrace{\left(\alpha_{s^{{}^{\prime}}}\frac{{\bm{x}}_{t}-\sigma_{t}\epsilon_{\bm{\theta}}({\bm{x}}_{t})}{\alpha_{t}}+\sigma_{s^{{}^{\prime}}}\epsilon_{\bm{\theta}}({\bm{x}}_{t})\right)}_{\text{predicted}\;{\bm{x}}_{s^{{}^{\prime}}}={\bm{f}}({\bm{x}}_{t},t,s^{{}^{\prime}})}+\underbrace{\sqrt{1-\frac{\alpha_{s}^{2}}{\alpha_{s^{{}^{\prime}}}^{2}}}{\bm{z}}}_{\text{controllable noise}},\end{split}$ (33)
+instead of $\eta$ in Equation 32, we introduced an stochastic parameter $\gamma$ that allows the control of the stochasticity in every step.
+As previously noted, the SSS framework is derived from the parameterisation of TCF and possesses the capability to explicitly manage stochastic intensities.
+Consequently, we define $\gamma$ as the parameterisation of target timestep $s$, with the linear range between full stochasticity and determinacy, thus $\gamma=1-s^{\prime}/s$ and $\eta=\sqrt{1-{\alpha_{s}^{2}}/{\alpha_{s^{{}^{\prime}}}^{2}}}$.
+Sampling Process.
+Similar to Multistep consistency sampling, we can simply decompose Equation 33 into the format of multi-step sampling, which is demonstrated in Algorithm 4:
+Denoise: ${\bm{x}}_{t\rightarrow s^{\prime}}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}^{\rightarrow s^{\prime}}({\bm{x}}_{t},t)$ (34)
+Diffuse: $\hat{{\bm{x}}}_{s}\leftarrow\frac{\alpha_{s}}{\alpha_{s^{\prime}}}{\bm{x}}_{t\rightarrow s^{\prime}}+\sqrt{1-\frac{\alpha_{s}^{2}}{\alpha_{s^{\prime}}^{2}}}{\bm{z}}$ (35)
 Specifically, every sampling step in SSS includes the denoise sub-step according to the ODE solver and the diffuse sub-step based on Langevin SDE.
 In comparison with multistep consistency sampling, where the endpoint and noise level are fixed, SSS introduces the additional parameter $\gamma$ to control the destination point for the denoise step and allows for the adjustment of random noise level for the diffuse step, as detailed in Figure 2(b) and Algorithm 4.
-This parameter is referred to as the stochastic parameter.
-In the denoising sub-step, our focus is on reducing the errors that arise from the $n$-th step in Equation 11 to $\mathcal{O}(\sqrt{\tau_{n}-(1-\gamma)\tau_{(n+1)}})$ when $n\in[\![1,N-1]\!]$, by the prediction of the non-origin, and ultimately optimize the accumulated error as shown in Theorem 4.2.
-Theorem 4.2.
-As the strategic stochastic sampling process defined in Algorithm 4, and denoting the $N>1$ steps sampling results with trained trajectory consistency model ${\bm{f}}_{\bm{\theta^{*}}}$ as $q_{\bm{\theta^{*}},N}={\bm{f}}_{{\bm{\theta^{*}}},\tau_{N}}\sharp\mathcal{N}(\alpha_{\tau_{N}}{\bm{x}}_{\tau_{(N-1)}\rightarrow 0},\sigma_{\tau_{N}}\bm{I})$, the results with exact trajectory consistency function ${\bm{f}}_{\bm{\theta}}$ as $p_{N}$, then the accumulated error in SSS is defined as:
-$TV(q_{\bm{\theta^{*}},N},p_{N})=\mathcal{O}\left(\sum_{n=1}^{N-1}\sqrt{\tau_{n}-(1-\gamma)\tau_{n+1}}+\sqrt{\tau_{N}}\right),$ (23)
-where $\gamma\in[0,1]$ is the parameter controlling the destination in each denoising step.
-Proof.
-The proof is provided in Section D.3. ∎
-While sufficient stochasticity in diffuse sub-step helps reduce both the estimation errors and discretisation accumulated in earlier sampling steps and drives the sample towards the desired marginal distribution, a similar finding is also observed in (Karras et al., 2022; Xu et al., 2023b).
-In SSS, we encourage the step size of the diffuse sub-step to be smaller than that of the denoise sub-step, which is contrary to (Xu et al., 2023b).
-It is also worth noting that when $\gamma$ is low, the estimation error plays a more important role.
+In (Kim et al., 2023), the authors proposed a unified sampling method and have delicately crafted a $\gamma$-sampler for its $G_{\theta}$ function to leverage its ability for “anytime to anytime” prediction.
+This sampler incorporates a $\gamma$ parameter to regulate the decoder $G_{\theta}$ target, yet further endows it with the properties of both EDM sampler and deterministic sampler within the Multistep Consistency Sampling protocol.
+Another key finding in (Kim et al., 2023) is the analysis of the error bound in $\gamma$-sampling $\mathcal{O}(\sum_{n=0}^{N-1}\sqrt{t_{n}-\sqrt{1-\gamma^{2}}t_{n+1}})$, which aligns with the previous practical observation (Karras et al., 2022): less stochastic will lead to a better sampling quality.
+It is worth noting that SSS was specially designed to control the stochasticity in LCM sampling with the well-defined $\gamma$ drived from DDIM sampling.
+Additionally, its alternating sampling framework is grounded on multistep consistency sampling, similar to the $\gamma$ sampling in (Kim et al., 2023).
+From this perspective, we notice that SSS is also capable of changing the stochastic intensity during its sampling procedure and has the same error bound as stated in (Kim et al., 2023).
+In another word, our method is also able to reduce the errors that arise from the $n$-th step to $\mathcal{O}(\sqrt{\tau_{n}-(1-\gamma)\tau_{(n+1)}})$ when $n\in[\![1,N-1]\!]$ and ultimately optimize the accumulated error.
+It is worth noting that in the process of sampling with TCF(1) and TCF(2), there is another discretization error $\mathcal{O}(h_{\Delta\tau}^{k+1})$ that cannot be eliminated, which is brought by the parameterisation of TCF as discussed in Equation 23.
+We would like to highlight that the parameterisation form of TCF makes it versatile for community samplers, and our SSS provides an additional method to explicitly control the stochastic level for LCM.
+It is also worth noting that when $\gamma$ is low, the estimation error plays a more important role (Lyu et al., 2023).
 Hence, the optimal value of $\gamma$ should be determined empirically, as we show in Section 5.3.
-Figure 4: Qualitative comparison. For each prompt, images are generated using the same random seed for every model, without any cherry-picking. More results are provided in Section E.3.
+Figure 4: Qualitative comparison. For each prompt, images are generated using the same random seed for every model, without any cherry-picking. More results are provided in Section D.3.
 Figure 5: Qualitative effects of stochastic parameter $\gamma$ with same NFEs.
@@
 The leftmost image is sampled from LCM (Luo et al., 2023b).
-More samples can be found in Section E.2.
+More samples can be found in Section D.2.
 Table 1: Quantitative comparison on the COCO validation set.
@@
 Figure 6: Quantitative ablation on different stochastic parameter $\gamma$.
+Table 2: Quantitative ablation on the TCF parameterisation type.
+Para Type FID $\downarrow$ IC Score $\uparrow$ ImageReward $\uparrow$ PickScore $\uparrow$
+TCF(1) 12.68 0.5095 68.49 22.31
+TCF(2) 13.35 0.5037 58.13 22.07
+TCF(S+) 13.03 0.4176 57.96 22.01
+Figure 7: Versatility of TCD.
+Our TCD LoRA can be directly applied to a wide range of models, including custom community models, styled LoRA, ControlNet, and IP-Adapter, accelerating their generation in just a few steps with high quality.
 4.4 Extension to Large Text Conditional Models
@@
 This results in the transformation of the trajectory function to ${\bm{f}}_{\rm{\theta}}({\bm{x}}_{t},{\bm{c}},t,s)$, and the guided distillation method proposed by (Meng et al., 2023; Luo et al., 2023a) can be directly integrated into it, as detailed in Algorithm 2.
-Trajectory consistency distillation can directly occur on top of the parameters of a pre-trained diffusion model as the fine-tuning process.
+Trajectory consistency distillation can be considered as fine-tuning process, which directly occur on top of the pre-trained model.
 For scaling TCD to larger models (e.g., SDXL) with significantly reduced memory consumption, we incorporate Low-Rank Adaptation (LoRA) (Hu et al., 2021), a parameter-efficient fine-tuning method, into the distillation process.
@@
 For detailed implementation information, please refer to Appendix C.
-Figure 7: Versatility of TCD.
-Our TCD LoRA can be directly applied to a wide range of models, including custom community models, styled LoRA, ControlNet, and IP-Adapter, accelerating their generation in just a few steps with high quality.
-Table 2: Quantitative ablation on the TCF parameterisation type.
-Para Type FID $\downarrow$ IC Score $\uparrow$ ImageReward $\uparrow$ PickScore $\uparrow$
-TCF(1) 12.68 0.5095 68.49 22.31
-TCF(2) 13.35 0.5037 58.13 22.07
-TCF(S+) 13.03 0.4176 57.96 22.01
 5.2 Main Results
@@
 We employed the Frechet Inception Distance (FID) and the Image Complexity Score (Feng et al., 2022) to assess the sample quality and visual complexity of generated images.
-Additionally, we used Image Reward (Xu et al., 2023a) and PickScore (Kirstain et al., 2023) to quantify text-image alignment and human preference.
+Additionally, we used Image Reward (Xu et al., 2023) and PickScore (Kirstain et al., 2023) to quantify text-image alignment and human preference.
 As depicted in Table 1, TCD shows an improvement in the performance of LCM across various timesteps and metrics.
@@
 Due to the introduction of additional parameters, TCF(S+) faces challenges in effectively distilling the teacher’s information, leading to poorer performance.
+Note that while TCF(1) only requires 3000 training iterations to achieve satisfactory results in Table 2, TCF(S+) takes 43,000 iterations but still lags behind in performance.
 Versatility of TCD.
@@
 The results shown in Figure 7 imply that TCD can be directly applied to various models to accelerate image generation with high quality in only 2-8 steps.
-Additional samples can be found in Section E.4.
+Additional samples can be found in Section D.4.
 6 Conclusion
-In this work, we introduce TCD, a novel distillation method that reduces inherent errors present in consistency models, including TCF for training and SSS for sampling.
-The TCF is proposed to diminish distillation errors and enable the model to track the trajectory along the PF ODE.
-Furthermore, SSS is proposed to reduce accumulated errors by the bijective traversal.
+In this work, we introduce TCD, a improved latent consistency distillation method that reduces inherent errors present in latent consistency models, including TCF for training and SSS for sampling.
+The TCF is proposed to diminish parameterisation and distillation errors and enable the model to track the trajectory along the PF ODE with a semi-linear structure.
+Furthermore, SSS is proposed to explicitly control the stochastic intensity and reduce accumulated errors by the bijective traversal.
 Remarkably, TCD outperforms LCM across all sampling steps and exhibits superior performance compared to numerical methods of teacher model.
@@
 Additionally, it is worth investigating an improved design to achieve fewer steps generation, e.g., single step.
-Impact Statements
-While our advancements in sample quality and speed can help reduce inference costs, they also have the potential to amplify negative societal effects, such as disseminating disinformation.
-In the future, it is encouraged to implement adequate guardrails and detection techniques to minimize the risk of misuse.
-Appendix A Related Works
-Diffusion Models.
-Diffusion Models (DMs), also known as Score-based generative models, have shown superior performance in various generative fields.
-Chen et al.
-(2022) provides theoretical convergence guarantees, implying that DMs can efficiently sample from essentially any realistic data distribution under minimal data assumptions.
-ADM (Dhariwal & Nichol, 2021) firstly shows the potential of DMs to outperform GANs.
-EDM (Karras et al., 2022) further elucidates the design space of GMs, clearly separating the concrete design choices, deriving best practices for the sampling process, and improving the training dynamics, thereby drastically improving the results.
-Text-conditional Diffusion Models.
-DMs have specially achieved great success in the area of text-to-image synthesis (Nichol et al., 2022; Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022; Balaji et al., 2022; Podell et al., 2023).
-To reduce computational cost, diffusion models typically operate within the latent space (Rombach et al., 2022; Podell et al., 2023) or include separate super-resolution steps (Ramesh et al., 2022; Saharia et al., 2022; Balaji et al., 2022).
-The integration of classifier-free diffusion guidance (Ho & Salimans, 2022; Dieleman, 2022) during the sampling process dramatically improves samples produced by conditional diffusion models at almost no additional cost.
-Fast Sampling of DMs.
-DMs exhibit great generative capabilities but are bottlenecked by their slow sampling speed.
-Various numerical methods driven by the SDE mathematical underpinnings underlying DMs have been proposed for accelerating DM sampling.
-DDIM (Song et al., 2020a) has originally shown promise for few-step sampling.
-Other works involve predictor-corrector samplers (Song et al., 2020b; Karras et al., 2022), exponential integrators (Lu et al., 2022a; Zhang & Chen, 2022; Lu et al., 2022b), and automated methods to adjust solvers (Kong & Ping, 2021; Jolicoeur-Martineau et al., 2021).
-Despite these methods having achieved great improvement in fast sampling, they are limited by the inherent discretisation error present in all solvers (De Bortoli et al., 2021).
-Ultimately, the sample quality obtained with few NFEs is limited.
-Another series of research works exemplified by distillation techniques (Luhman & Luhman, 2021; Salimans & Ho, 2022; Meng et al., 2023).
-These methods distill knowledge from pretrained models into few-step samplers, representing an efficient solution for few NFEs.
-However, they may experience a lengthy and costly process that requires huge amounts of data and suffer from slow convergence.
-Consistency Models.
-To overcome the limitations of current fast samplers, Song et al.
-(2023) proposes Consistency Models (CMs) learning a direct mapping from noise to data built on the top of the trajectory of PF ODE, achieving one-step generation while allowing multi-step sampling to trade compute for quality.
-Lyu et al.
-(2023) provide the first convergence guarantees for CMs under moderate assumptions.
-Kim et al.
-(2023) proposes a universal framework for CMs and DMs.
-The core design is similar to ours, with the main differences being that we focus on reducing error in CMs, subtly leverage the semi-linear structure of the PF ODE for parameterization, and avoid the need for adversarial training.
-Latent Consistency Models (LCMs) (Luo et al., 2023a) integrate consistency distillation with latent diffusion models (Rombach et al., 2022) to achieve impressive performance in accelerating text-to-image synthesis.
-LCM-LoRA (Luo et al., 2023b) further improves the training efficiency and versatility of LCMs by introducing LoRA (Hu et al., 2021) into the distillation process.
-However, due to inherent accumulated errors in multistep consistency samplings, LCM experiences a notable decline in quality at high NFEs.
-In contrast, our TCD leverages trajectory consistency and bijective traversal to eliminate this defect.
-Parameter-Efficient Fine-Tuning.
-Training a diffusion model is highly resource-intensive and environmentally unfriendly.
-Fine-tuning such a model can also be challenging due to the vast number of parameters involved (Aghajanyan et al., 2020).
-Thus, Parameter-Efficient Fine-Tuning (PEFT) (Houlsby et al., 2019) was proposed to enable the fine-tuning of pretrained models with a limited number of parameters required for training.
-Among these techniques, Low-Rank Adaptation (LoRA) (Hu et al., 2021) has demonstrated transcendental performance.
-LoRA’s strategy involves freezing the pretrained model weights and injecting trainable rank decomposition matrices, which succinctly represent the required adjustments in the model’s weights for fine-tuning.
-With this strategy, LoRA significantly reduces the volume of parameters to be modified, thereby substantially decreasing both computational load and storage demands.
+Appendix A Discussion between CTM (Kim et al., 2023) and TCD
+A recent work, CTM (Kim et al., 2023), proposed a unified framework for CM and DM solvers, with the idea of “anytime to anytime” prediction, theoretically substantiated and facilitated through a $G_{\theta}$ function with $\gamma$-sampling for assessing both the integrand and the integral of PF-ODE.
+Conversely, TCD focused on improving the performance of LCM by leveraging a semi-linear structure with a narrower time spacing in the consistency function.
+While these two studies share intersections in their approaches, such as the use of varying time intervals from $t$ to $s$, they also exhibit distinct differences.
+In this section, we delve into a discussion on CTM and TCD, exploring their various aspects.
+To initiate our comparison, we first recall two definitions: explicit numerical method and implicit numerical method for ODEs (Atkinson et al., 2009; Süli, 2010):
+A general explicit method can be:
+$y_{n+1}=y_{n}+hf(y_{n},t_{n}),\,\,n=0,\dots,N-1,$ (36)
+where $f(y_{n},t_{n})$ is a continuous function and $h$ is the interval between $t_{n}$ and $t_{n+1}$. In an explicit method, $y_{n}$ is already calculated, and $y_{n+1}$ can be given directly.
+Similarly, the implicit method, which is also said to be absolutely stable, can be expressed in the form:
+$y_{n+1}=y_{n}+hf(y_{n+1},t_{n+1}),\,\,n=0,\dots,N-1,$ (37)
+which requires the solution of an implicit equation $f(y_{n+1},t_{n+1})$ in order to determine $y_{n+1}$.
+A representative example of an explicit method is the Euler Method, while the Backward Euler Method is an implicit method.
+Informally, score-based sampling with time-discretized numerical integral solvers can be regarded as an explicit method, whereas the distillation model that directly estimates the integral of ODEs is an implicit method (Kim et al., 2023).
+In order to unify CMs and DMs by assessing both the infinitesimal jump and long jump, CTM (Kim et al., 2023) proposed a unified decoder $G$ with an “anytime-to-anytime” idea, where $G$ is the solution of the PF-ODE with an integrator term.
+Furthermore, CTM used a trainable $g_{\theta}$ to parameterize the integrator term and made it access the infinitesimal jump inspired by the Euler Solver.
+In order to train such a $G_{\theta}$ function, instead of directly forcing $G_{\theta}({\bm{x}}_{t},t,s)\approx\texttt{Solver}({\bm{x}}_{t},t,s;\phi)$, CTM proposed a soft consistency matching $G_{\theta}({\bm{x}}_{t},t,s)\approx G_{\text{sg}(\theta)}(\texttt{Solver}({\bm{x}}_{t},t,u;\phi),u,s)$ that supports both $u=s$ and $u=t-\Delta t$.
+It can be observed that this loss is similar to the distillation loss for consistency model (Song et al., 2020a).
+Additionally, CTM has also introduced DSM loss and GAN loss to optimize the final performance of the model.
+It’s evident that the adversarial loss plays an important role in reaching a better FID.
+To further leverage the implicit method with a well-trained $g_{\theta}$ function, CTM proposed $\gamma$-sampling, which integrates the properties of both the EDM sampler and the deterministic ODE sampler, and falls within the protocol of multistep consistency sampling, i.e., iterative denoising and diffusing.
+If we were to apply the informal analogy of explicit and implicit methods, CTM proposed a distillation-based implicit method with $G_{\theta}$ to address the problems in time-discretized explicit solvers.
+In contrast, TCD aims to tackle the challenge posed by the imprecise parameterisation of the consistency function (Song et al., 2023) in LCM (Luo et al., 2023a) and improve the generation quality on text-to-image task.
+Our method strictly adheres to the definitions of the consistency model, enabling the direct adoption of characteristics and algorithms inherent to consistency models.
+We first observed that several errors in the distillation process are related to the time interval $t\rightarrow s$ of the consistency function.
+Thus, we leverage the semi-linear structure with exponential integrators of the PF-ODE for parameterization, which also supports a shorter interval (i.e., moderately moving the upper limit $s$ away from 00).
+Such concept of transitioning from $t$ to $s$ overlaps with CTM’s “anytime to anytime” idea but ours is from different view.
+Moreover, this semi-linear structure supports ${\bm{x}}_{0}$, ${\bm{v}}$, and $\epsilon$ parameterisation, allowing for efficient distillation from large-scale diffusion models.
+During training, we simply follow all the definitions and targets of Consistency Distillation and equip $\mathbb{E}[\omega(t_{n},t_{m})\|{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n+k}},t_{n+k},t_{m})-{\bm{f}}_{\bm{\theta}^{-}}(\hat{{\bm{x}}}_{t_{n}}^{\phi,k},t_{n},t_{m})\|^{2}_{2}]$ as the loss while $\hat{{\bm{x}}}_{t_{n}}^{\phi,k}$ is from a solver by 1 NFE with parameter $\phi$ and $k\geq 1$.
+As observed in (Sauer et al., 2023), GAN loss tends to result in poorer mode coverage, hence we only employ the distillation loss used in (Song et al., 2023; Luo et al., 2023a) to ensure the diversity of generated images.
+The sampling process of TCD remains explicit as the TCF can be easily parameterized by $\epsilon$, allowing it to directly utilize existing explicit samplers.
+This means that once trained, TCF functions as a plug-and-play model, eliminating the need for designing additional samplers.
+Recognizing Multistep Consistency Sampling with parameterisation consistency function in LCM as a fixed stochasticity DDIM solver (Equation 31), we introduced SSS to achieve a flexible stochastic intensity control.
+Since CTM (Kim et al., 2023) proposed a unified framework for CMs and DMs and our SSS is directly based on multistep sampling in CMs, it demonstrates theoretical similarity to the $\gamma$-sampling used in CTM (Kim et al., 2023).
+And the theoretical insights provided by CTM (Kim et al., 2023) in sampling errors can also support our conjectures, e.g., injecting more noise will lead to degraded quality.
+In summary, CTM has four key contributions: 1.
+A single model inspired by Euler Solver that attains all the advantages of score-based models and distillation models ; 2.
+$\gamma$-sampling for flexible sampling; 3.
+Soft matching for efficient and stabilized training; 4.
+The use of auxiliary losses for better student training.
+Meanwhile, TCD is an enhancement of LCM (Luo et al., 2023a), adhering to the definition of consistency models with the following features: 1.
+Better parameterization of consistency functions with a semi-linear structure (Lu et al., 2022a, b; Zhang & Chen, 2022); 2.
+Maintains the explicit sampling method and is versatile for existing solvers; 3.
+Employs only consistency loss with skipping-steps to achieve self-consistency and avoid mode collapse.
+Moreover, TCD aims to improve the quality of general generation tasks in fewer steps, instead of achieving SOTA for single-step generation for specific tasks.
 Appendix B Algorithmic Details
 B.1 Algorithm Details of Trajectory Consistency Distillation
-Due to space limitations, we omitted some implementation details in the main body, but we provided a detailed algorithm for trajectory consistency distillation during training as Algorithm 1 and a guided version for conditional models as Algorithm 2.
+We provided a detailed algorithm for trajectory consistency distillation during training as Algorithm 1 and a guided version for conditional models as Algorithm 2.
 Algorithm 1 Trajectory Consistency Distillation
@@
      Sample ${\bm{x}}_{t_{n+k}}\sim\mathcal{N}(\alpha_{t_{n+k}}{\bm{x}},\sigma_{t_{n+k}}^{2}\bm{I})$
-     $\hat{{\bm{x}}}_{t_{n}}^{\phi,k}\leftarrow(1+\omega)\Phi^{(k)}({\bm{x}}_{t_{n+k}},t_{n+k},t_{n},{\bm{c}};\bm{\phi})-$ $\omega\Phi^{(k)}({\bm{x}}_{t_{n+k}},t_{n+k},t_{n},{\bm{c}};\bm{\phi})$
+     $\hat{{\bm{x}}}_{t_{n}}^{\phi,k}\leftarrow(1+\omega)\Phi^{(k)}({\bm{x}}_{t_{n+k}},t_{n+k},t_{n},{\bm{c}};\bm{\phi})-$ $\omega\Phi^{(k)}({\bm{x}}_{t_{n+k}},t_{n+k},t_{n},\varnothing;\bm{\phi})$
      $\mathcal{L}(\bm{\theta},\bm{\theta^{-}};\bm{\phi})\leftarrow$ $\|{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n+k}},{\bm{c}},t_{n+k},t_{m})-{\bm{f}}_{\bm{\theta}^{-}}(\hat{{\bm{x}}}_{t_{n}}^{\phi,k},{\bm{c}},t_{n},t_{m})\|^{2}_{2}$
@@
      ${\bm{x}}_{\tau_{n}\rightarrow\tau_{n}^{{}^{\prime}}}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}(\tilde{{\bm{x}}}_{\tau_{n}},\tau_{n})$ Diffuse
-     $\tilde{{\bm{x}}}_{\tau_{(n+1)}}\leftarrow\alpha_{\tau_{(n+1)}}{\bm{x}}_{\tau_{n}\rightarrow\tau_{n}^{{}^{\prime}}}+\sigma_{\tau_{(n+1)}}{\bm{z}}$
+     $\hat{{\bm{x}}}_{\tau_{(n+1)}}\leftarrow\alpha_{\tau_{(n+1)}}{\bm{x}}_{\tau_{n}\rightarrow\tau_{n}^{{}^{\prime}}}+\sigma_{\tau_{(n+1)}}{\bm{z}}$
   end for
@@
      Calculate ending time $\tau_{n}^{{}^{\prime}}=(1-\gamma)\tau_{(n+1)}$ Denoise
-     ${\bm{x}}_{\tau_{n}\rightarrow\tau_{n}^{{}^{\prime}}}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}(\tilde{{\bm{x}}}_{\tau_{n}},\tau_{n},\tau_{n}^{{}^{\prime}})$ Diffuse
+     ${\bm{x}}_{\tau_{n}\rightarrow\tau_{n}^{{}^{\prime}}}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}^{\rightarrow\tau_{n}^{{}^{\prime}}}(\hat{{\bm{x}}}_{\tau_{n}},\tau_{n})$ Diffuse
      $\tilde{{\bm{x}}}_{\tau_{(n+1)}}\leftarrow\frac{\alpha_{\tau_{(n+1)}}}{\alpha_{\tau_{n}^{{}^{\prime}}}}{\bm{x}}_{\tau_{n}\rightarrow\tau_{n}^{{}^{\prime}}}+\sqrt{1-\frac{\alpha_{\tau_{(n+1)}}^{2}}{\alpha_{\tau_{n}^{{}^{\prime}}}^{2}}}{\bm{z}}$
   end for
-  ${\bm{x}}_{\tau_{N}\rightarrow 0}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}(\tilde{{\bm{x}}}_{\tau_{N}},\tau_{N},0)$
+  ${\bm{x}}_{\tau_{N}\rightarrow 0}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}^{\rightarrow 0}(\tilde{{\bm{x}}}_{\tau_{N}},\tau_{N})$
   Output: ${\bm{x}}_{\tau_{N}\rightarrow 0}$
@@
 The inference time in Figure 1 was tested on an A800, averaged over 16 prompts.
-Appendix D Proofs
-D.1 Proof of Corollary 3.1
-Before we prove the corollary 3.1, we assume the following mild conditions on the data distribution $p_{\rm{data}}$ (Lee et al., 2023; Chen et al., 2022):
-Assumption D.1.
-The data distribution has finite second moment, that is, $\mathbb{E}_{{\bm{x}}_{0}\sim p_{\rm{data}}}[\|{\bm{x}}_{0}\|_{2}^{2}]=m_{2}^{2}<\infty$.
-Assumption D.2.
-The score function $\nabla\log p_{t}({\bm{x}})$ is Lipschitz on the variable ${\bm{x}}$ with Lipschitz constant $L_{s}\geq 1$, $\forall t\in[0,T]$.
-We further assume bounds on the score estimation error and consistency error (Song et al., 2023; Lyu et al., 2023):
-Assumption D.3.
-Assume $\mathbb{E}_{{\bm{x}}_{t_{n}}\sim p_{t_{n}}}[\|{\bm{s}}_{\bm{\phi}}({\bm{x}}_{t_{n}},t_{n})-\nabla\log p_{t_{n}}({\bm{x}}_{t_{n}})\|_{2}^{2}]\leq\varepsilon_{\text{sc}}^{2},\forall n\in[\![1,N]\!]$.
-Assumption D.4.
-Assume $\mathbb{E}_{{\bm{x}}_{t_{n}}\sim p_{t_{n}}}[\|{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n+1}},t_{n+1})-{\bm{f}}_{{\bm{\theta}}}(\hat{\bm{x}}_{t_{n}}^{\bm{\phi}},t_{n})\|_{2}^{2}]\leq\varepsilon_{\text{cm}}^{2}(t_{n+1}-t_{n})^{2},\forall n\in[\![1,N-1]\!]$
-Assumption D.5.
-The consistency model ${\bm{f}}_{\bm{\theta}}({\bm{x}},t_{n})$ is Lipschitz on the variable ${\bm{x}}$ with Lipschitz constant $L_{f}>1,~{}\forall n\in[\![1,N]\!].$
-We also follow the (Lyu et al., 2023) to assume a discrete schedule:
-Assumption D.6.
-Assume the discretisation schedule $0<\delta=t_{1}<t_{2}<\cdots<t_{N}=T$, $h_{k}=t_{k+1}-t_{k}$ to Equation 3 is divided into two stages:
-  1. 1.
-    $h_{k}\equiv h$ for all $k\in[\![N_{1},N-1]\!],$ and $(N-N_{1}-1)h<T\leq(N-N_{1})h$;
-  2. 2.
-    $h_{k}=2^{-(N_{1}-k)}h=\frac{h_{k+1}}{2}$ for $k\in[\![1,N_{1}-1]\!]$, $N_{1}$ satisfies $h_{2}=2^{-(N_{1}-2)}h\leq 2\delta.$
-Then we introduce some properties and a lemma borrowed from (Lyu et al., 2023), which shows that TV error can be bounded after a small time OU regularisation.
-The forward OU process can be defined as follows:
-$\mathrm{d}{\bm{x}}_{t}=-{\bm{x}}_{t}\mathrm{d}t+\sqrt{2}\mathrm{d}{\bm{w}}_{t},$ (26)
-and denote the Markov kernel $P_{\text{OU}}^{s}$ to be defined by Equation 26, that is, if ${\bm{x}}_{t}\sim p$ for some $p$ be a distribution over $\mathbb{R}^{d}$, then ${\bm{x}}_{t+s}\sim pP_{\text{OU}}^{s}$.
-Under Assumptions D.1-D.6, when $T>L_{s}^{-1}$, the one-step generating error is bounded as follows:
-$W_{2}({\bm{f}}_{{\bm{\theta}},T}\sharp\mathcal{N}({\bm{0}},{\bm{I}}_{d}),p_{\rm{data}})\lesssim(d^{\frac{1}{2}}\vee m_{2})L_{f}e^{-T}+T(\varepsilon_{\text{cm}}+L_{f}\varepsilon_{\text{sc}}+L_{f}L_{s}^{\frac{3}{2}}d^{\frac{1}{2}}h)+(d^{\frac{1}{2}}\vee m_{2})\delta^{\frac{1}{2}}.$ (27)
-Lemma D.7.
-For any two distributions $p$ and $q$, running the OU process in Equation 26 for $p,q$ individually with time $\tau>0$, the following TV distance bound holds,
-[…]
-Proof.
-Denote $\psi_{\sigma^{2}}({\bm{y}})$ as the density function to the normal distribution $\mathcal{N}({\bm{0}},\sigma^{2}{\bm{I}}_{d})$.
-We write the $\text{TV}(pP_{\text{OU}}^{\tau},qP_{\text{OU}}^{\tau})$ into integral form as:
-$\text{TV}(pP_{\text{OU}}^{\tau},qP_{\text{OU}}^{\tau})$ $=\frac{1}{2}\int_{\mathbb{R}^{d}}|(pP_{\text{OU}}^{\tau})({\bm{x}})-(qP_{\text{OU}}^{\tau})({\bm{x}})|\mathrm{d}{\bm{x}}$
-$=\frac{1}{2}\int_{\mathbb{R}^{d}}\left|\int_{\mathbb{R}^{d}}p({\bm{y}})\psi_{1-e^{-2\tau}}({\bm{x}}-e^{-\tau}{\bm{y}})\mathrm{d}{\bm{y}}-\int_{\mathbb{R}^{d}}q({\bm{z}})\psi_{1-e^{-2\tau}}({\bm{x}}-e^{-\tau}{\bm{z}})\mathrm{d}{\bm{z}}\right|\mathrm{d}{\bm{x}}.$ (28)
-Take a coupling $\gamma\in\Gamma(p,q)$, then
-$\int_{\mathbb{R}^{d}}\gamma({\bm{y}},{\bm{z}})\mathrm{d}{\bm{z}}=p({\bm{y}}),$
-$\int_{\mathbb{R}^{d}}\gamma({\bm{y}},{\bm{z}})\mathrm{d}{\bm{y}}=q({\bm{z}}),$ (29)
-we have
-$\text{TV}(pP_{\text{OU}}^{\tau},qP_{\text{OU}}^{\tau})$ $=\frac{1}{2}\int_{\mathbb{R}^{d}}\left|\int_{\mathbb{R}^{\mathrm{d}\times d}}\gamma({\bm{y}},{\bm{z}})[\psi_{1-e^{-2\tau}}({\bm{x}}-e^{-\tau}{\bm{y}})-\psi_{1-e^{-2\tau}}({\bm{x}}-e^{-\tau}{\bm{z}})]\mathrm{d}{\bm{y}}\mathrm{d}{\bm{z}}\right|\mathrm{d}{\bm{x}}.$
-$\leq\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d\times d}}\gamma({\bm{y}},{\bm{z}})\left|\psi_{1-e^{-2\tau}}({\bm{x}}-e^{-\tau}{\bm{y}})-\psi_{1-e^{-2\tau}}({\bm{x}}-e^{-\tau}{\bm{z}})\right|\mathrm{d}{\bm{y}}\mathrm{d}{\bm{z}}\mathrm{d}{\bm{x}}$
-$=\int_{\mathbb{R}^{d\times d}}\gamma({\bm{y}},{\bm{z}})\left(\frac{1}{2}\int_{\mathbb{R}^{d}}\left|\psi_{1-e^{-2\tau}}({\bm{x}}-e^{-\tau}{\bm{y}})-\psi_{1-e^{-2\tau}}({\bm{x}}-e^{-\tau}{\bm{z}})\right|\mathrm{d}{\bm{x}}\right)\mathrm{d}{\bm{y}}\mathrm{d}{\bm{z}}$
-$=\int_{\mathbb{R}^{d\times d}}\gamma({\bm{y}},{\bm{z}})\text{TV}(\psi_{1-e^{-2\tau}}(\cdot-e^{-\tau}{\bm{y}}),\psi_{1-e^{-2\tau}}(\cdot-e^{-\tau}{\bm{z}}))\mathrm{d}{\bm{y}}\mathrm{d}{\bm{z}}$
-$\leq\int_{\mathbb{R}^{d\times d}}\gamma({\bm{y}},{\bm{z}})\sqrt{\frac{1}{2}\text{KL}(\psi_{1-e^{-2\tau}}(\cdot-e^{-\tau}{\bm{y}})\|\psi_{1-e^{-2\tau}}(\cdot-e^{-\tau}{\bm{z}}))}\mathrm{d}{\bm{y}}\mathrm{d}{\bm{z}}$
-$=\int_{\mathbb{R}^{d\times d}}\gamma({\bm{y}},{\bm{z}})\frac{1}{2}\sqrt{\frac{e^{-2\tau}}{1-e^{-2\tau}}\|{\bm{y}}-{\bm{z}}\|_{2}^{2}}\mathrm{d}{\bm{y}}\mathrm{d}{\bm{z}}$
-$=\frac{1}{2\sqrt{e^{2\tau}-1}}\int_{\mathbb{R}^{d\times d}}\gamma({\bm{y}},{\bm{z}})\|{\bm{y}}-{\bm{z}}\|_{2}\mathrm{d}{\bm{y}}\mathrm{d}{\bm{z}}$ (30)
-Noting $\frac{1}{2\sqrt{e^{2\tau}-1}}\leq\frac{1}{2\sqrt{2\tau}}$, and taking $\gamma$ over all coupling $\Gamma(p,q)$, we have
-$\text{TV}(pP_{\text{OU}}^{\tau},qP_{\text{OU}}^{\tau})\lesssim\frac{1}{\sqrt{\tau}}W_{1}(p,q)\leq\frac{1}{\sqrt{\tau}}W_{2}(p,q)$
-Let $q$ be the output of consistency models, either the one step consistency sampling result or the $k$-th multistep consistency sampling result.
-To control the TV error, we smooth the generated sample by the forward OU process with a small time that is the same as the early stopping time $\delta$, and then we can get the TV distance between $q_{k}P_{\text{OU}}^{\delta}$ and $p_{\rm{data}}$.
-Proof of Corollary 3.1.
-According to the triangular inequality, Lemma D.7 and Equation 27,
-$\text{TV}(q_{1}P_{\text{OU}}^{\delta},p_{\rm{data}})$ $\leq\text{TV}(q_{1}P_{\text{OU}}^{\delta},p_{\delta}P_{\text{OU}}^{\delta})+\text{TV}(p_{\delta}P_{\text{OU}}^{\delta},p_{\rm{data}})$
-$\lesssim\frac{1}{\sqrt{\delta}}W_{2}(q_{1},p_{\delta})+\text{TV}(p_{2\delta},p_{\rm{data}})$
-$\lesssim\frac{1}{\sqrt{\delta}}\left((d^{\frac{1}{2}}\vee m_{2})L_{f}e^{-T}+T(\varepsilon_{\text{cm}}+L_{f}\varepsilon_{\text{sc}}+L_{f}L_{s}^{\frac{3}{2}}d^{\frac{1}{2}}h)\right)+\text{TV}(p_{2\delta},p_{\rm{data}}).$ (31)
-Note that if we take $\delta\asymp\frac{\varepsilon^{2}}{L_{s}^{2}(d\vee m_{2}^{2})}$, then by Lemma 6.4, (Lee et al., 2023), $\text{TV}(p_{2\delta},p_{\rm{data}})\leq\varepsilon$, this concludes that
-$\text{TV}(q_{1}P_{\text{OU}}^{\delta},p_{\rm{data}})\lesssim\frac{L_{s}(d^{\frac{1}{2}}\vee m_{2})}{\varepsilon}[(\log L_{f}+\frac{T}{2^{k}})(\varepsilon_{\text{cm}}+L_{f}\varepsilon_{\text{sc}}+L_{f}L_{s}^{\frac{3}{2}}d^{\frac{1}{2}}h)+\frac{(d^{\frac{1}{2}}\vee m_{2})L_{f}}{2^{k}e^{T}}]+\varepsilon.$
-Similarly for $q_{k}$, if we take $\delta\asymp\frac{\varepsilon^{2}}{L_{s}^{2}(d\vee m_{2}^{2})}$
-$\text{TV}(q_{k}P_{\text{OU}}^{\delta},p_{\rm{data}})$ (32)
-$\leq$ $\text{TV}(q_{k}P_{\text{OU}}^{\delta},p_{\delta}P_{\text{OU}}^{\delta})+\text{TV}(p_{\delta}P_{\text{OU}}^{\delta},p_{\rm{data}})$
-$\lesssim$ $\frac{L_{s}(d^{\frac{1}{2}}\vee m_{2})}{\varepsilon}[(\log L_{f}+\frac{T}{2^{k}})(\varepsilon_{\text{cm}}+L_{f}\varepsilon_{\text{sc}}+L_{f}L_{s}^{\frac{3}{2}}d^{\frac{1}{2}}h)+\frac{(d^{\frac{1}{2}}\vee m_{2})L_{f}}{2^{k}e^{T}}]+\varepsilon.$ (33)
-D.2 Proof of Theorem 4.1
-The proof closely follows Song et al. (2023) and relies on induction, drawing parallels with the classic proof of global error bounds for numerical ODE solvers (Prasad, 1990).
-Proof.
-For simplicity of notation, we denote $\bm{\theta^{*}}$ as $\bm{\theta}$. From $\mathcal{L}_{\rm{TCD}}^{N}(\bm{\theta},\bm{\theta};\bm{\phi})=0$, we have
-$\mathcal{L}^{N}_{\rm{TCD}}=\mathbb{E}\left[\omega(t_{n},t_{m})\left\|{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n+1}},t_{n+1},t_{m})-{\bm{f}}_{\bm{\theta}}(\hat{{\bm{x}}}_{t_{n}}^{\bm{\phi}},t_{n},t_{m})\right\|^{2}_{2}\right]=0.$ (34)
-According to the definition, it follows that $p_{t_{n}}({\bm{x}}_{t_{n}})>0$ for every $x_{t_{n}}$ and $1\leqslant n\leqslant N$. Therefore,  Equation 34 entails
-$\omega(t_{n},t_{m})\left\|{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n+1}},t_{n+1},t_{m})-{\bm{f}}_{\bm{\theta}}(\hat{{\bm{x}}}_{t_{n}}^{\bm{\phi}},t_{n},t_{m})\right\|^{2}_{2}\equiv 0.$ (35)
-Because $\lambda(\cdot)>0$ and $\left\|{\bm{x}},{\mathbf{y}}\right\|^{2}_{2}=0\Leftrightarrow{\bm{x}}={\mathbf{y}}$, this further implies that
-${\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n+1}},t_{n+1},t_{m})\equiv{\bm{f}}_{\bm{\theta}}(\hat{{\bm{x}}}_{t_{n}}^{\bm{\phi}},t_{n},t_{m}).$ (36)
-Let ${\bm{e}}_{n}^{s}$ represent the error vector that arises when predicting $x_{s}$ at $t_{n}$, which is defined as
-${\bm{e}}_{n,m}:={\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n}},t_{n},t_{m})-{\bm{f}}({\bm{x}}_{t_{n}},t_{n},t_{m};\bm{\phi}).$
-We can easily derive the following recursion relation
-${\bm{e}}_{n+1,m}$ $={\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n+1}},t_{n+1},t_{m})-{\bm{f}}({\bm{x}}_{t_{n+1}},t_{n+1},t_{m};\bm{\phi})$ (37)
-$={\bm{f}}_{\bm{\theta}}(\hat{{\bm{x}}}_{t_{n}}^{\bm{\phi}},t_{n},t_{m})-{\bm{f}}({\bm{x}}_{t_{n}},t_{n},t_{m};\bm{\phi})$
-$={\bm{f}}_{\bm{\theta}}(\hat{{\bm{x}}}_{t_{n}}^{\bm{\phi}},t_{n},t_{m})-{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n}},t_{n},t_{m})+{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n}},t_{n},t_{m})-{\bm{f}}({\bm{x}}_{t_{n}},t_{n},t_{m};\bm{\phi})$
-$={\bm{f}}_{\bm{\theta}}(\hat{{\bm{x}}}_{t_{n}}^{\bm{\phi}},t_{n},t_{m})-{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{n}},t_{n},t_{m})+{\bm{e}}_{n,m}.$
-Because ${\bm{f}}_{\theta}(\cdot,t_{n},t_{m})$ has Lipschitz constant $L$, we have
-$\lVert{\bm{e}}_{n+1,m}\rVert_{2}$ $\leqslant\lVert{\bm{e}}_{n,m}\rVert_{2}+L\left\lVert\hat{{\bm{x}}}_{t_{n}}^{\bm{\phi}}-{\bm{x}}_{t_{n}}\right\rVert_{2}$
-$=\lVert{\bm{e}}_{n,m}\rVert_{2}+L\cdot\mathcal{O}((t_{n+1}-t_{n})^{p+1})$
-$=\lVert{\bm{e}}_{n,m}\rVert_{2}+\mathcal{O}((t_{n+1}-t_{n})^{p+1}.$
-In addition, we observe that ${\bm{e}}_{m,m}=\bm{0}$, because
-${\bm{e}}_{m,m}$ $={\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{m}},t_{m},t_{m})-{\bm{f}}({\bm{x}}_{t_{m}},t_{m},t_{m};\bm{\phi})$
-=$\overset{(i)}{=}{\bm{x}}_{t_{m}}-{\bm{f}}({\bm{x}}_{t_{m}},t_{m},t_{m};\bm{\phi})$
-=$\overset{(i)}{=}{\bm{x}}_{t_{m}}-{\bm{x}}_{t_{m}}$
-$=\bm{0},$
-here $(i)$ is true because trajectory consistency distillation restricts the boundary condition for the parameterized model such that ${\bm{f}}_{\bm{\theta}}({\bm{x}}_{t_{m}},t_{m},t_{m})={\bm{x}}_{t_{m}}$, and $(ii)$ is entailed by the definition of trajectory function ${\bm{f}}(\cdot,\cdot,\cdot;\bm{\phi})$ in  Equation 13.
-This allows us to perform induction on the recursion formula Equation 37 to obtain
-$\lVert{\bm{e}}_{n,m}\rVert_{2}$ $\leqslant\lVert{\bm{e}}_{m,m}\rVert_{2}+\sum_{k=m}^{n-1}\mathcal{O}((t_{k+1}-t_{k})^{p+1}$
-$=\sum_{k=m}^{n-1}\mathcal{O}((t_{k+1}-t_{k})^{p+1}$
-$=\sum_{k=m}^{n-1}(t_{k+1}-t_{k})\mathcal{O}((t_{k+1}-t_{k})^{p}$
-$\leqslant\sum_{k=m}^{n-1}(t_{k+1}-t_{k})\mathcal{O}((\Delta t)^{p}$
-$=\mathcal{O}((\Delta t)^{p}\sum_{k=m}^{n-1}(t_{k+1}-t_{k})$
-$=\mathcal{O}((\Delta t)^{p}(t_{n}-t_{m}),$
-which completes the proof. ∎
-D.3 Proof of  Theorem 4.2
-In this section, our derivation mainly borrows the proof from (Kim et al., 2023; Chen et al., 2022).
-For DDPM (Chen et al., 2022),
-Theorem D.8.
-Suppose that Assumptions D.2D.1, and D.3 hold.
-Let $q_{T}$ be the output of the DDPM algorithm at time $T$, and suppose that the step size $h:=T/N$ satisfies $h\lesssim 1/L_{s}$, where $L_{s}\geq 1$ is the Lipschitz constant of score function.
-Then, it holds that
-[…]
-For simplicity, we assume the optimal TCD, $f_{\bm{\theta}^{*}}\equiv f$ with a well-learned $\bm{\theta}^{*}$, which recovers the true $f(\cdot,t,s)$ function.
-We establish that the density propagated by this optimal TCD model from any time $t$ to a subsequent time $s$ aligns with the predefined density determined by the fixed forward process.
-We now present the proposition ensuring the alignment of the transited density.
-Proposition D.9.
-[…]
-  1. 1.
-    Linear growth: $\|\nabla\log p_{t}(\mathbf{x})\|_{2}\leq L(t)(1+\|\mathbf{x}\|_{2})$, for all $\mathbf{x}\in\mathbb{R}^{D}$
-  2. 2.
-    Lipschitz: $\|\nabla\log p_{t}(\mathbf{x})-\nabla\log p_{t}(\mathbf{y})\|_{2}\leq L(t)\|\mathbf{x}-\mathbf{y}\|_{2}$, for all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{D}$.
-Then for any $t\in[0,T]$ and $s\in[0,t]$, $p_{s}=\mathcal{T}_{t\rightarrow s}\sharp p_{t}$.
-This theorem guarantees that by learning the optimal TCD, which possesses complete trajectory information, we can retrieve all true densities at any time using TCD.
-Define $\mathcal{T}_{t\rightarrow s}$ as the oracle transition mapping from $t$ to $s$ via the diffusion process in Equation 1.
-Let $\mathcal{T}_{t\rightarrow s}^{\bm{\theta}^{*}}(\cdot)$ represent the transition mapping from the optimal TCD model, and $\mathcal{T}_{t\rightarrow s}^{\bm{\phi}}(\cdot)$ represent the transition mapping from the empirical probability flow ODE.
-Since all processes start at point $T$ with an initial probability distribution $p_{T}$ and $\mathcal{T}_{t\rightarrow s}^{\bm{\theta}^{*}}(\cdot)=\mathcal{T}_{t\rightarrow s}^{\bm{\phi}}(\cdot)$Theorem D.8 and $\mathcal{T}_{T\rightarrow t}\sharp p_{T}=p_{t}$ from Proposition D.9 tell us that for $t>s$
-$\text{TV}\left(\mathcal{T}_{t\rightarrow s}\sharp p_{t},\mathcal{T}_{t\rightarrow s}^{\bm{\theta}^{*}}\sharp p_{t}\right)=\text{TV}\left(\mathcal{T}_{t\rightarrow s}\sharp p_{t},\mathcal{T}_{t\rightarrow s}^{\bm{\phi}}\sharp p_{t}\right)=\mathcal{O}(\sqrt{t-s}).$ (38)
-$\text{TV}\left(\mathcal{T}_{t\rightarrow 0}\mathcal{T}_{(1-\gamma)t\rightarrow t}\mathcal{T}_{T\rightarrow(1-\gamma)t}\sharp p_{T},\mathcal{T}_{t\rightarrow 0}^{\bm{\theta}^{*}}\mathcal{T}_{(1-\gamma)t\rightarrow t}\mathcal{T}_{T\rightarrow(1-\gamma)t}^{\bm{\theta}^{*}}\sharp p_{T}\right)$
-$\overset{(a)}{\leq}$ $\text{TV}\left(\mathcal{T}_{t\rightarrow 0}\mathcal{T}_{(1-\gamma)t\rightarrow t}\mathcal{T}_{T\rightarrow(1-\gamma)t}\sharp p_{T},\mathcal{T}_{t\rightarrow 0}^{\bm{\theta}^{*}}\mathcal{T}_{(1-\gamma)t\rightarrow t}\mathcal{T}_{T\rightarrow(1-\gamma)t}\sharp p_{T}\right)$
-$+$ $\text{TV}\left(\mathcal{T}_{t\rightarrow 0}^{\bm{\theta}^{*}}\mathcal{T}_{(1-\gamma)t\rightarrow t}\mathcal{T}_{T\rightarrow(1-\gamma)t}\sharp p_{T},\mathcal{T}_{t\rightarrow 0}^{\bm{\theta}^{*}}\mathcal{T}_{(1-\gamma)t\rightarrow t}\mathcal{T}_{T\rightarrow(1-\gamma)t}^{\bm{\theta}^{*}}\sharp p_{T}\right)$
-=$\overset{(b)}{=}$ $\text{TV}\left(\mathcal{T}_{t\rightarrow 0}\mathcal{T}_{T\rightarrow t}\sharp p_{T},\mathcal{T}_{t\rightarrow 0}^{\bm{\theta}^{*}}\mathcal{T}_{T\rightarrow t}\sharp p_{T}\right)+\text{TV}\left(\mathcal{T}_{T\rightarrow(1-\gamma)t}\sharp p_{T},\mathcal{T}_{T\rightarrow(1-\gamma)t}^{\bm{\theta}^{*}}\sharp p_{T}\right)$
-=$\overset{(c)}{=}$ $\text{TV}\left(\mathcal{T}_{t\rightarrow 0}\sharp p_{t},\mathcal{T}_{t\rightarrow 0}^{\bm{\theta}^{*}}\sharp p_{t}\right)+\text{TV}\left(\mathcal{T}_{T\rightarrow(1-\gamma)t}\sharp p_{T},\mathcal{T}_{T\rightarrow(1-\gamma)t}^{\bm{\theta}^{*}}\sharp p_{T}\right)$
-=$\overset{(d)}{=}$ $\mathcal{O}(\sqrt{t})+\mathcal{O}(\sqrt{T-(1-\gamma)t}).$
-Here (a) is obtained from the triangular inequality, (b) and (c) are due to $\mathcal{T}_{(1-\gamma)t\rightarrow t}\mathcal{T}_{T\rightarrow(1-\gamma)t}=\mathcal{T}_{T\rightarrow t}$ and $\mathcal{T}_{T\rightarrow t}\sharp p_{T}=p_{t}$ from Proposition D.9, and (d) comes from Equation 38.
-Appendix E Additional Results
-E.1 More Samples with Different NFEs
+Appendix D Additional Results
+D.1 More Samples with Different NFEs
 In Figure 8, we present additional samples synthesized by LCM and TCD with varying NFEs.
 These samples consistently demonstrate that the details of images generated by LCM tend to vanish as the NFEs increase, whereas TCD either maintains or enhances image details at higher NFEs.
-E.2 More Samples with Different Stochastic Parameter $\gamma$
+D.2 More Samples with Different Stochastic Parameter $\gamma$
 In Figure 9, we showcase additional samples generated by TCD with varying stochastic parameter $\gamma$, demonstrating that the visual complexity and quality of the image gradually improve as $\gamma$ increases.
-E.3 More Comparisons Results
+D.3 More Comparisons Results
 In Figure 10 to Figure 14, we present additional examples for comparison with state-of-the-art methods.
@@
 Additionally, it can produce more detailed and high-definition images with sufficient NFEs, even surpassing the performance of DPM-Solver++(2S).
-E.4 More samples from TCD’s Versatility Testing
+D.4 More Samples from TCD’s Versatility Testing
 To ascertain the versatility of TCD, we test it on a wide range of models, including the popular community model Animagine XL V3  [^7] The checkpoint of Animagine can be found in https://civitai.com/models/260267/animagine-xl-v3, LoRA Papercut XL  [^8] The checkpoint of Papercut can be found in https://civitai.com/models/122567/papercut-sdxl, Depth ControlNet  [^9] Depth ControlNet: https://huggingface.co/diffusers/controlnet-depth-sdxl-1.0, Canny ControlNet  [^10] Canny ControlNet: https://huggingface.co/diffusers/controlnet-canny-sdxl-1.0, and IP-Adapter  [^11] IP-Adapter: https://github.com/tencent-ailab/IP-Adapter.